### v1

# FIDLAR: Forecast-Informed Deep Learning Architecture for Flood Mitigation

Abstract
In coastal river systems, frequent floods, often occurring during major storms or king tides, pose a severe threat to lives and property. However, these floods can be mitigated or even prevented by strategically releasing water before extreme weather events with hydraulic structures such as dams, gates, pumps, and reservoirs. A standard approach used by local water management agencies is the “rule-based” method, which specifies predetermined pre-releases of water based on historical and time-tested human experience, but which tends to result in excess or inadequate water release. The model predictive control (MPC), a physics-based model for prediction, is an alternative approach, albeit involving computationally intensive calculations. In this paper, we propose a Forecast Informed Deep Learning Architecture, FIDLAr, to achieve rapid and optimal flood management with precise water pre-releases. FIDLAr seamlessly integrates two neural network modules: one called the Flood Manager, which is responsible for generating water pre-release schedules, and another called the Flood Evaluator, which assesses these generated schedules. The Evaluator module is pre-trained separately, and its gradient-based feedback is used to train the Manager model, ensuring optimal water pre-releases. We have conducted experiments using FIDLAr with data from a flood-prone coastal area in South Florida, particularly susceptible to frequent storms. Results show that FIDLAr is several orders of magnitude faster than currently used physics-based approaches while outperforming baseline methods with improved water pre-release schedules. Our code is at https://github.com/JimengShi/FIDLAR/.
1 Introduction
Floods can result in catastrophic consequences with considerable loss of life Jonkman and Vrijling (2008), huge socio-economic impact Wu et al. (2021), property damage Brody et al. (2007), and environmental devastation Yin et al. (2023a). They pose a threat to food and water security as well as sustainable development Kabir and Hossen (2019). What is even more alarming is that research indicates global climate change may lead to a drastic increase in flood risks in terms of both frequency and scale Wing et al. (2022); Hirabayashi et al. (2013). In coastal areas, flood risks are increasing due to sea-level rise Sadler et al. (2020). Therefore, effective flood management is of utmost importance.
To mitigate flood risks, water management agencies have built controllable hydraulic structures such as dams, gates, pumps, and reservoirs in river systems Kerkez et al. (2016). Determining optimal control schedules of these hydraulic structures is however a challenging problem Bowes et al. (2021). Rule-based methods Bowes et al. (2021); Sadler et al. (2019) formulate control schedules based on insights gained from historically observed data. These schedules are often formalized and extensively documented in databases Marchese et al. (2018). They represent the collective wisdom gathered over decades of experience in managing specific river systems and are commonly adopted by local water management authorities. However, these schedules may expose vulnerabilities during extreme events, which are becoming more likely due to climate change, and might lack sufficient locally available historical data, particularly for events like 500- or 1000-year floods. Moreover, these schedules may not offer effective solutions for complex river systems Schwanenberg et al. (2015).
Model Predictive Control (MPC) is a well-known strategy that tackles this challenge as an optimization process Karimanzira (2016). By optimizing variables such as water levels and discharges, the algorithm generates control schedules for hydraulic structures to effectively mitigate overall flood risks Delaney et al. (2020); Zarei et al. (2021). Random initialization followed by soft-computing techniques such as genetic algorithms is often employed to perform the optimizations. Subsequently, physics-based models such as HEC-RAS and SWMM are used to assess the generated control schedules. Publications on this topic include the following Leon et al. (2020); Sadler et al. (2019); Vermuyten et al. (2018); Chen et al. (2016); Leon et al. (2014); Yin et al. (2023b). However, such MPC methods are prohibitively slow since they require thousands of simulations using physics-based models.
In the last decade or so, Machine learning (ML) has emerged as a powerful tool in this domain. So far, the ML-based methods have been applied to flood prediction Mosavi et al. (2018); Shi et al. (2023b), flood detection Tanim et al. (2022); Shahabi et al. (2020), susceptibility assessment Saha et al. (2021); Islam et al. (2021), and post-flood management Munawar et al. (2019). However, ML-based methods have not been applied to flood mitigation tasks. We aim to fill this gap. Moreover, we replace the compute-intensive MPC approach with a framework based purely on deep learning (DL), where both the generator and evaluator are achieved by DL modules.
Gradient-based planning with world models aims to seek optimal actions by training a world model and conducting inferences using that trained model. We refer the reader to Figure 1(a) from Jyothir et al. (2023).
Figure 1: Forecast-Informed Deep Learning Architecture (FIDLAr). Input data in the form of a table consisting of five variables as shown in the left end of the diagram. The variables $w$ and $k$ are the lengths of the past and prediction windows, respectively. The parts of the table colored green are provided as inputs to the programs, while the orange and blue parts (with question marks) are outputs. The black and red arrows represent the propagation and back-propagation processes, respectively. The Flood Manager and Flood Evaluator represent DL models, the former to predict the control schedules of gates and pumps to pre-release water, and the latter to determine water levels for those control schedules. Loss functions, $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, penalize the flooding and water wastage beyond pre-specified thresholds, respectively.
Our proposed Forecast Informed Deep Learning Architecture, FIDLAr, uses gradient-based planning to achieve optimal flood management with rapid and feasible water pre-releasing. See Figure 1 for details on its architecture. FIDLAr’s characteristics are summarized as follows:
  • FIDLAr seamlessly combines two DL models in series: Flood Manager and Flood Evaluator. The former model is responsible for generating water pre-release schedules, while the latter model accurately forecasts the resulting water levels. The latter model serves as an evaluator for the output of the former. During training, the gradient feedback from the Evaluator model can be back-propagated to reinforce the Manager model to generate effective water pre-release schedules. The Manager and Evaluator play the role of Planner and World Model, respectively, from the model from Jyothir et al. (2023).
  • FIDLAr is a data-driven approach, learning flood mitigation strategies from historically observed data. Once trained, it offers rapid response capabilities. This highlights the advantages of DL-based models over physics-based models, particularly in the context of real-time flood management.
  • FIDLAr is a model-agnostic framework, where both the Manager and Evaluator could be any type of DL model with differentiable loss functions that allow back-propagation.
  • FIDLAr is trained with a custom loss function to balance both flood risks and water wastage at the same time, which is a multitask learning method.
  • FIDLAr makes effective use of reliable forecasts of specific variables (e.g., precipitation and tidal information) for the near future.
  • The explainability features of FIDLAr allow us to better understand the behavior of complex river systems.
2 Problem Formulation
Flood mitigation involves the prediction of control schedules for hydraulic structures like gates and pumps within the river system, denoted as $X^{gate,pump}_{t+1:t+k}$, spanning $k$ time points from $t+1$ to $t+k$ into the future. This prediction uses as input historical input data (on all variables), $X$, from the preceding $w$ time points, in conjunction with reliably forecasted covariates (such as rainfall and tide) for the next $k$ time points. Thus, we train a DL model whose transfer function, $\mathcal{F}_{\theta}$, has the form:
$\mathcal{F}_{\theta}:(X_{t-w+1:t},X^{cov}_{t+1:t+k})\rightarrow X^{gate,pump}_{t+1:t+k},$ (1)
where the subscripts represent the time ranges, and the superscripts refer to the variables under consideration. Superscripts are dropped when all variables are taken into account. The superscript $cov$ refers specifically to the reliably predicted covariates (e.g., rain, tides).
3 Methodology
3.1 The Challenge
Intuitively, an ML model can be trained to learn the function $f_{\theta}$. However, the challenge here is that the historical data may not represent the desired values for the control schedules of the gates and pumps, especially if it resulted in flood conditions, and therefore cannot be used as the ground-truth values used in traditional supervised learning. Therefore, we separate a “planner” (Flood Manager) module from an “evaluator” (Flood Evaluator) module. We first build an independent and accurate evaluator with the historical data, which attempts to learn the consequences of every possible action. We then treat the planner module design as an optimization problem where we optimize a custom loss function, allowing the system to pick the action that would prevent undesirable consequences. Desired states of the river systems avoid floods on one hand while preventing water wastage on the other. Once the Flood Evaluator is trained, it is frozen and used to train the Flood Manager. The evaluator serves to backpropagate the gradient feedback regarding the quality of the schedules generated by the manager, thus guiding the manager to generate more effective schedules of gates and pumps. Such a training process ends up refining the function, $f_{\theta}$, as represented in Eq. (1).
Figure 2: Flood Evaluator. Unlike Figure 1, the parts shaded green (includes historical data and covariates predicted from near future) and orange (representing any control schedule for the gates and pumps) are both used as inputs. Water levels (blue) are the outputs.
3.2 Flood Evaluator
Flood Evaluator is tasked with accurately forecasting water levels at designated points of interest within river systems for any control schedule of gates and pumps provided to it. See Figure 2. The underlying transfer function of the Evaluator has the form:
$\mathcal{G}_{\theta}:(X_{t-w+1:t},X^{cov}_{t+1:t+k},X^{gate,pump}_{t+1:t+k})\rightarrow X^{water}_{t+1:t+k}.$ (2)
As mentioned earlier, the Evaluator module is trained independently using the historical data to achieve highly accurate water level predictions for any given set of conditions and control schedules. Therefore, once the Evaluator is trained, its parameters are set in stone for the training of the Manager, where it plays the role of a trained “referee”.
Figure 3: Flood Manager. As in Figure 1, the parts shaded green (historical data) are the inputs, and the parts shaded orange are the outputs. The water levels shaded blue are not predicted.
3.3 Flood Manager
As shown in Figure 3, the Flood Manager is trained to produce optimal control schedules for the hydraulic structures (gates and pumps), taking reliably predictable future information (rain, tide) and all historical data. During its training, since no ground truth is available, it is trained using the gradient descent algorithm Ruder (2016); Andrychowicz et al. (2016), using the partial derivatives of the loss function (Eq. (3)), which in turn needs the evaluator.
3.4 Custom Loss Function
Loss functions are critical in steering the learning process. For flood management, since no ground truth is available, independent metrics of performance are needed. An obvious one is the total time (Figure 4 (a)) for which the water levels either exceed the flooding threshold or dip below the water wastage threshold. Another related metric is the extent to which the limits are exceeded to signify the severity of floods or water wastage (Figure 4 (b)). The lower threshold for flood management is important in practice, since it prevents water wastage, thereby supporting irrigation, facilitating navigation, and maintaining ecological balance. It also prevents the optimization methods from trivially recommending the depletion of valuable water resources to prevent future flooding. The above metrics suggest the following components for the loss function for our models. The final loss function is a balanced combination of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ in Eq. (4).
$\mathcal{L}_{1}=\sum^{N}_{i=1}\sum^{t+k}_{j=t+1}\Arrowvert max\{\hat{X}^{water}_{i,j}-X^{flood}_{i},0\}\Arrowvert^{2},$ (3)
$\mathcal{L}_{2}=\sum^{N}_{i=1}\sum^{t+k}_{j=t+1}\Arrowvert min\{\hat{X}^{water}_{i,j}-X^{waste}_{i},0\}\Arrowvert^{2},$
where $N$ is the number of water level locations of interest; $k$ is the length of prediction horizon; $X^{flood}$ and $X^{waste}$ represent the thresholds for flooding and water wastage; and their capped versions, $\hat{X}^{flood}$ and $\hat{X}^{waste}$ are obtained using the evaluator module. The combined loss function is given by:
$\mathcal{L}_{total}=\alpha\cdot\mathcal{L}_{1}+\beta\cdot\mathcal{L}_{2},$ (4)
where $\alpha/\beta$ dictates the relative importance of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$.
Figure 4: The two red bars represent an upper threshold of flooding and a lower threshold of wastage. Shown are (a) the time spans when these thresholds are crossed, and (b) the areas between the water level curve and the threshold bars. Violations of the upper (lower, resp.) threshold are captured in $\mathcal{L}_{1}$ ($\mathcal{L}_{2}$, resp.).
Figure 5: Architecture of Graph Transformer Network for Flood Evaluator. Input variables include Rainfall, Pump, Gate, Tide, and Water levels as shown in Figure 6 and Table 1, which are generically denoted by $R,T,G,P$, and $W$. Output is water levels. We use the same architecture for Evaluator and Manager, but with minor changes accordingly of the input and output(see Figs. 2 and 3).
3.5 Training FIDLAr
The entire training process is outlined below. The Evaluator model is trained in a traditional way using the observed historical data as the ground truth for predictions and minimizing a loss function that penalizes deviations from the ground truth. After the Evaluator model is trained, its parameters are frozen and the Manager model is trained. For every control schedule that the Manager generates, $X^{gate,pump}_{t+1:t+k}$, the Evaluator is used to compute the resulting water levels, which is then used to compute the loss function defined in Eq. (4). Gradient descent Ruder (2016) provides the information to be back-propagated to be used to update the parameters of Manager. Thus,
${X^{*}}^{gate,pump}_{t+1:t+k}=\underset{X^{gate,pump}_{t+1:t+k}}{\arg\min}\{\mathcal{L}_{total}\}.\$ (5)
The training details of FIDLAr are given in Algorithm 1.
4 GTN Model
The Manager and Evaluator modules described so far are model agnostic. We tried many existing architectures for them, as discussed in Section 5.2. The Graph Transformer Network (GTN) architecture combines graph neural networks (GNNs), attention-based transformer networks, long short-term memory networks (LSTMs), and the convolutional neural networks (CNNs). GNN and LSTM modules are combined for learning the spatiotemporal dynamics of water levels, while the Transformer and CNN modules focus on extracting feature representations from the covariates. The Attention mechanism is employed to discern the interactions between covariates and water levels, as shown in Eq. (6) below. Figure 5 presents the GTN architecture used (with minor modifications) for the Evaluator and the Manager.
$Atte(Q,K,V){}$ $=softmax(\frac{Q^{cov}(K^{water})^{T}}{\sqrt{d_{k}}})V^{water}$ (6)
$=softmax(\frac{Q^{water}(K^{cov})^{T}}{\sqrt{d_{k}}})V^{cov},$
where $T$ denotes the transpose operation; $water$ and $cov$ represent the variables of water levels and covariates.
Algorithm 1 Training algorithm of FIDLAr
Input: recent past data: ${\bf X}_{t-w+1,t}$
Input: near future data: ${\bf X}^{cov}_{t+1,t+k}={\bf X}^{rain,tide,gate,pump}_{t+1,t+k}$
Parameter: $\theta_{E},\theta_{M}$: parameters of Evaluator and Manager
Parameter: $w,k:$ length of past and prediction windows
1:  // Train Flood Evaluator, $\mathcal{G}_{\theta_{E}}$
2:  initialize learnable parameters $\theta_{E}$
3:  for $i=1,\ldots,N$ epochs do
4:     $\text{MiniBatch}\leftarrow(\{{\bf X}_{t-w+1,t},{\bf X}^{cov}_{t+1,t+k}\},{\bf X}^{water}_{t+1,t+k})$
5:     $\hat{{\bf X}}^{water}_{t+1,t+k}\leftarrow\mathcal{G}_{\theta_{E}}({\bf X}_{t-w+1,t},{\bf X}^{cov}_{t+1,t+k})$
6:     $\mathcal{L}_{E}\leftarrow\frac{1}{k}\sum_{j=1}^{k}||\hat{{\bf X}}^{water}_{j}-{\bf X}^{water}_{j}||^{2}$
7:     $\nabla_{\theta_{E}}\leftarrow\text{BackwardAD}(\mathcal{L}_{E})$
8:     $\theta_{E}\leftarrow\theta_{E}-\eta\nabla_{\theta_{E}}$
9:  end for
10:  return trained Flood Evaluator, $\mathcal{G}_{\theta_{E}}$
11:  // Train Flood Manager, $\mathcal{F}_{\theta_{M}}$, with frozen $\mathcal{G}_{\theta_{E}}$
12:  initialize learnable parameters $\theta_{M}$
13:  while $X^{water}_{t,t+k}$ violates either threshold do
14:     $\text{MiniBatch}\leftarrow(\{{\bf X}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k}\},{\bf X}^{gate,pump}_{t+1,t+k})$
15:     $\hat{{\bf X}}^{gate,pump}_{t+1,t+k}\leftarrow\mathcal{F}_{\theta_{M}}({\bf X}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k})$
16:     $\hat{{\bf X}}^{water}_{t+1,t+k}\leftarrow\mathcal{G}_{\theta_{E}}({\bf X}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k},\hat{{\bf X}}^{gate,pump}_{t+1,t+k})$
17:     $\mathcal{L}_{E}=\alpha\cdot\mathcal{L}_{1}(\hat{{\bf X}}^{water}_{t+1,t+k})+\beta\cdot\mathcal{L}_{2}(\hat{{\bf X}}^{water}_{t+1,t+k})$
18:     $\nabla_{\theta_{M}}\leftarrow\text{BackwardAD}(\mathcal{L}_{E})$
19:     $\theta_{M}\leftarrow\theta_{M}-\eta\nabla_{\theta_{M}}$
20:  end while
21:  return trained Flood Manager, $\mathcal{F}_{\theta_{M}}$
5 Experiments
5.1 Data
We obtained data from the South Florida Water Management District’s (SFWMD) DBHydro database District (2023) for the coastal stretch of the South Florida watershed. The data set consists of hourly observations for water levels and external covariates from Jan. 1st, 2010 to Dec. 31, 2020, a period totaling eleven years. As shown in Figure 6, the river system under consideration has two branches and includes several hydraulic structures (gates, pumps) to control water flows. Water levels in coastal systems are also impacted by tides. We aim to predict effective schedules on hydraulic structures (gates, pumps) to minimize flood risks at four specific locations marked by green circles in Figure 6. To highlight the risk posed by floods, we note that this portion of the river system flows through Miami downtown, with a sizable population, many commercial enterprises, and an international airport in its vicinity.
Figure 6: Schemetic diagram of study domain - South Florida. There are three water stations with hydraulic structures (S26, S25B, S25A), one simple water station, S1 (green circle in the middle), and one station monitoring tide information from the ocean.
Table 1: Summary of the data set (Jan. 1, 2010 $\sim$ Dec. 31, 2020).
Feature Interval Unit #Var. Location
Rainfall Hourly $inch/h$ 1 -
Tide Hourly $ft$ 1 S4
Pump Hourly $ft^{3}/s$ 2 S25B, S26
Gate Hourly $ft$ 3 S25A, S25B, S26
Water Hourly $ft$ 4 S25A, S25B, S26, S1
5.2 Experimental designs
The sliding input window Li et al. (2014) (also known as look-back window Gidea and Katz (2018) strategy was used to process the entire dataset Shi et al. (2023a)), which helps to ensure that the input and output vector lengths are constant. The datasets were split in chronological order with the first 80% for training and the remaining 20% for testing.
Flood Prediction.
The primary role of the Flood Evaluator is to forecast flood events by predicting water levels for given input conditions. To assess its performance, we compared several DL models for the evaluation, including our proposed GTN model as well as a physics-based model (HEC-RAS). We measured accuracy using multiple metrics: (a) the mean absolute error (MAE), (b) the root mean square error (RMSE) computed between the predicted and actual water levels, (c) the number of time points where the upper or lower thresholds were breached, and (d) the area between the water level curve and the threshold bars. For consistency, we used a look-back window of length $w=72$ hours and a prediction window of length $k=24$ hours. The eight DL methods used are listed below (more details in Appendix B):
  • MLP Suykens et al. (1995): Multilayer perceptron can learn complex non-linear dependencies;
  • RNN Medsker and Jain (2001): Recurrent neural networks are good at processing sequential data;
  • CNN O’Shea and Nash (2015): A 1D convolutional neural network;
  • GNN Kipf and Welling (2016): Graph neural network with nodes representing variables and edges representing spatial dependencies;
  • TCN Bai et al. (2018): Temporal dilated convolutional network with an exponentially large receptive field;
  • RCNN Zhang and Dong (2020): Combined RNN and CNN model for time series forecasting;
  • Transformer Vaswani et al. (2017): Attention-based network for sequence modeling (only encoder);
  • GTN: Combining GNNs with LSTMs, CNNs, and transformers, as described in Figure 5.
Flood Mitigation with FIDLAr.
FIDLAr requires both Evaluator and Manager components and we chose the GTN model for the Evaluator, since it was the best performing model. For the Manager model, we experimented with one rule-based method, and two genetic algorithms – one with a physics-based HEC-RAS evaluator Leon et al. (2020) and one with our DL-based GTN evaluator, and several DL-based managers. As shown in Table 3, the list of DL approaches used in our experiments for the Manager included MLP, RNN, CNN, GNN, TCN, RCNN, Transformer, and GTN. Performance of FIDLAr was measured using (a) the number of time steps where the upper/lower thresholds were exceeded for the water levels, and (b) the area between the water level curve and the threshold bars.
5.3 Analysis of computational time
Since FIDLAr was designed for real-time flood control, we measured the running times of the models used in this work.
Table 2: Comparison of model performances for the Flood Evaluator on the test set, specifically at time t+1 for measurement station S1. The terms “Over Timesteps” and “Under Timesteps” indicate the number of time steps during which water levels exceed the upper threshold or fall below the lower threshold, respectively. Similarly, “Over Area” and “Under Area” pertain to the area between the water level curve and upper or lower threshold, as was illustrated in Figure 4. Results in red are lowest in that column or closest to the ground truth (in blue).
Methods MAE (ft) RMSE (ft) Over Timesteps Over Area Under Timesteps Under Area
Ground-truth - - 96 14.82 1,346 385.80
HEC-RAS 0.174 0.222 68 10.07 1,133 325.33
MLP 0.065 0.086 147 27.96 1,677 500.41
RNN 0.054 0.072 110 17.12 1,527 441.41
CNN 0.079 0.104 58 5.91 1,491 413.22
GNN 0.054 0.070 102 15.90 1,569 462.63
TCN 0.050 0.065 47 5.14 1,607 453.63
RCNN 0.092 0.110 37 4.61 1,829 553.20
Transformer 0.050 0.066 151 25.95 1,513 434.13
GTN (ours) 0.040 0.056 100 15.64 1,390 398.84
Table 3: Comparison of model performances for the Flood Manager on the test set, specifically at time t+1 for measurement station S1. The $*{}^{*}$ denotes that the GA method was used with a physics-based (HEC-RAS) evaluator. The -- denotes that the experiments were timed out. The ${}^{{\dagger}}$ denotes that the GA method was used with the GTN as the evaluator. All other rows are DL-based flood managers with a DL-based GTN as the evaluator. Results in red font are the best in that column.
Method Manager Over Timesteps Over Area Under Timesteps Under Area
Rule-based 96 14.82 1,346 385.8
GA-based Genetic $*{}^{*}$ (GA) - - - -
Genetic ${}^{{\dagger}}$ (GA) 86 16.54 454 104
DL-based MLP 91 13.31 1,071 268.35
RNN 35 3.97 351 61.05
CNN 81 11.22 1,163 314.37
GNN 31 3.72 429 84.31
TCN 39 3.77 306 55.12
RCNN 29 3.28 328 58.68
Transformer 85 11.54 1,180 310.16
GTN (Ours) 22 2.23 299 53.34
6 Experimental Results
6.1 Flood Prediction
To evaluate the accuracy of the Flood Evaluator, we test FIDLAr with various DL models on the test set by predicting the water levels $k=24$ hours in advance. Table 2 shows the MAEs and RMSEs for the DL-based models and the physics-based model (HEC-RAS). As discussed earlier, we also analyze the time steps and areas over and under the thresholds. We set the upper threshold (flood level) at 3.5 feet and the lower threshold (wastage level) at 0.0 feet. However, the methods remain consistent for many reasonable choices of threshold values. From Table 2, we note that GTN surpasses the performance of other models with predictions most closely aligned with the ground truth (highlighted in red). More results are provided in Table 8 in Appendix C.1.
6.2 Flood Mitigation
Performance measures.
Table 3 shows that all DL-based methods consistently performed better for site S1 than rule-based and GA-based approaches. Moreover, GTN has the best performance under all four metrics, whether it is to control floods or water wastage. The results for all other measurement sites follow a similar pattern, as shown in Table 9 in Appendix C.2.
Performance visualization.
We visualize the mitigated water levels for a short sample spanning 18 hours from September 3rd (09:00) to September 4th (03:00) in 2019 for one location of interest. Figure 7 indicates that FIDLAr equipped with GTN has resulted in water levels within the upper and lower thresholds. The zoomed portion shows a 2.5-hour period where the water levels are decreased under 3.5 feet based on the predicted gate and pump schedules. We provide the corresponding performance in Table 10. More visuals are in Appendix D.
Figure 7: Visualization of water levels with various methods for flood mitigation. We zoomed in the period of $t=2\sim 4.5$ in gray.
6.3 Ablation study
The ablation study quantifies the contribution of each of the components of GTN, by measuring the performance of GTN after removing each of the individual components. Results are shown in Table 4.
Table 4: Ablation study for flood mitigation for the entire test set (for time point t+1 at S1). The last row indicates the performance of the FIDLAr system with GTN as proposed in Figure 5. The best results are in red font.
Removed Over Over Under Under
Method Timesteps Area Timesteps Area
CNN 37 4.37 476 85.54
Transformer 32 3.57 325 57.42
GNN 56 5.90 479 86.22
LSTM 35 4.34 329 56.74
Attention 32 3.59 341 60.48
GTN 22 2.23 299 53.34
6.4 Computational time
Table 5 shows the running times for all the methods for the evaluator component and for the whole flood mitigation system in its test phase. All the DL-based approaches are several orders of magnitude faster than the currently used physics-based and GA-based approaches. The table also shows the training times for the DL-based approaches, although they are not necessary for real-time performance.
Table 5: The running time for flood prediction and mitigation. The running time for the rule-based method is not reported since historical data was directly used. $*{}^{*}$, which combines a GA-based tool and HEC-RAS, took too long and was not reported although the time for a small data set is reported in Tables 10, 11 in Appendix C.3. ${}^{{\dagger}}$, which combines the GA-based tool with GTN, also took too long but was estimated using a smaller sample.
Model Prediction Mitigation
Train Test Train Test
HEC-RAS - 45 min - -
Rule-based - - - -
${}^{{\dagger}}$ est. 30 h
MLP 35 min 1.88 s 58 min 6.13 s
RNN 243 min 8.57 s 54 min 12.75 s
CNN 37 min 1.93 s 17 min 5.84 s
GNN 64 min 3.13 s 29 min 7.26 s
TCN 60 min 4.57 s 45 min 9.06 s
RCNN 136 min 8.61 s 61 min 13.27 s
Transformer 43 min 2.38 s 23 min 6.76 s
GTN 119 min 2.95 s 35 min 4.90 s
6.5 Model explainability
Attention-based methods allow us to calculate the “attention scores” assigned to an input variable to compute a specific output variable. This is exemplified in the heatmap shown in Figure 8, which shows the attention scores assigned to the tide (columns) to compute the gate schedule output (rows) for 24 hours into the future. Note that there are 96 columns and 24 rows because we use 72 hours of past tidal observations and 24 hours of future predicted tidal data to predict 24 hours of the gate schedule into the future. Thus the rows $[0,23]$ correspond to the 24 hours into the future, while the columns $[0,95]$ also include 72 hours of the recent past. Thus, $t=72$ corresponds to the “current” time point, and the columns $[72,95]$ correspond to the same time points as rows $[0,23]$.
Figure 8: Attention scores between tide input and one of the gate schedule outputs. $x$ and $y$ axes are the tide and gate over time.
7 Discussion
Interpreting Attention Map.
The explainability feature, which was shown with an example in Figure 8, can provide significant insights into our results. Firstly, the simple example with the attention paid to the variable “tide” shows that the brightest patches in the heatmap occur in the last 24 columns. This highlights the importance of the novel idea of using tidal forecasting information in our flood mitigation approach. Given that tidal information within a 24-hour window can be quite accurate, this is useful. While tides may have a more predictable pattern over time, the contribution of rain on the water levels can also be seen for other time points. A second critical insight is that the brightest attention patches are in columns where the tide is at its highest is critical to the prediction of gate schedules. Additionally, the water level at the first high tide peak after the “current” time is more significant than the other two. Third, the gate schedule has peaks at times $t=11h$ and $t=22h$ into the future, which correspond to the lowest points of the tide. This implies that the optimal time for pre-releasing water is during low tide phases. Opening gates during high tide periods in coastal river systems is less advisable, as it may lead to water flowing back upstream from the ocean. Finally, we observe that while there is a light patch around column $t=65h$, suggesting mild attention for the previous high tide peak, but almost no attention to any of the peaks prior to that. This again suggests that we could have chosen to use a smaller window for the past input. Doing this analysis for more time points could provide evidence for the right value of $w$, the size of the look-back window.
8 Conclusions
FIDLAr is a DL-based tool to compute water “pre-release” schedules for hydraulic structures in a river system to achieve effective and efficient flood mitigation, while ensuring that water wastage is avoided. This was made possible by the use of well-crafted loss functions for the DL models. The dual component design (with a Manager and an Evaluator) is a strength of FIDLAr. During training, the gradient-based back-propagation from the Evaluator helps to fine-tune the Manager. The reliability of forecasts of variables like precipitation and tides is well known, but their critical importance may have been best demonstrated for the first time using the attention mechanism employed herein.
All the DL-based versions of FIDLAr are several orders of magnitude faster than the (physics-based or GA-based) competitors while achieving improvement over other methods in flood mitigation. These characteristics allow us to entertain the possibility of real-time flood management, which was previously unthinkable.
Acknowledgments
This work is part of the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) project, which is funded by the National Science Foundation under award number 2118329.
Appendix A Data and Processing
Data.
[…]
Processing.
Flood Evaluator is used to predict the water levels given the input of past $w$ time steps and some information of future $k$ time steps, either predicted information (“Rain” and “Tide”) or pre-determined information (“Gate” and “Pump”). It serves as an evaluator to assess the gate and pump schedules by predicting the resulting water levels. Flood Manager is used to predict the gate and pump schedules given the input of past $w$ time steps and some information of future $k$ time steps, either predicted information (“Rain” and “Tide”). Both models are trained as supervised learning tasks. Therefore, we pre-processed the data into pairs consisting of the input and output described above. In our experiments, we set $w=72$ hours and $k=24$ hours. The shape of the input and output is $(96,N_{input})$ and $(24,N_{output})$, where $N_{input}$ and $N_{output}$ are the number of input and output variables, respectively. Among the input, the future data with question marks is masked to a $mask\_value=1e-10$ using Keras API [^1] https://keras.io/api/layers/core˙layers/masking/. If all values in the input tensor at that time steps are equal to $mask\_value$, then the time steps will be masked (skipped) in all downstream layers.
(a) Flood Evaluator
(b) Flood Manager
Figure 9: Flood Evaluator and Flood Manager in FIDLAR.
Appendix B Model Architecture and Training Hyper-parameter
B.1 Architecture
Since FIDLAR is a model-agnostic framework, we use different deep learning models below as the backbone. In this section, we provide the model architecture, accordingly.
  • MLP: We stacked two Dense layers with $64$ and $32$ neurons. Each Dense layer uses RELU as the activation function and $L_{1}$ and $L_{2}$ as the regularization factor. Each Dense layer is followed by a Dropout layer to avoid possible overfitting.
  • RNN: We used one SimpleRNN layer with $64$ neurons followed by a Dropout layer. The SimpleRNN layer uses RELU as the activation function and $L_{1}$ and $L_{2}$ as the regularization factor.
  • CNN: We stacked three CNN1D layers with $64$, $32$, and $16$ filters, followed by MaxPooling1D layer with pool size $=2$. Each CNN1D layer uses RELU as the activation function, “same” padding mode sand 2 as the kernel size.
  • GNN: We consider each time series as one node and the connection as the edge. For the variables in Fig. 6, we assign “1” to the edge between two connected nodes, otherwise, “0”. For the architecture, we stacked two graph convolutional networks (GCNs) with $32$ and $16$ neurons. In parallel, a LSTM layer with $32$ units is used to learn the temporal dynamics from the input. We then concatenate the embedding from GCN and LSTM and output the outcomes.
  • TCN: It is a temporal dilated convolutional network with an exponentially large receptive field. We used $64$ as the number of filters, $2$ as the kernel size, $[1,2,4,8,16,32]$ as dilations, and 0.1 as the dropout rate. Subsequently, we added two Dense layers to compute the output. LayerNormalization is employed after each layer.
  • RCNN: A SimpleRNN with $64$ neurons, RELU as the activation function and $L_{1}$ and $L_{2}$ as the regularization factor is used first. We then used a CNN1D layer with $32$ filters to deal with the embedding above. The CNN1D layer has 2 as kernel size, “same” padding mode, and RELU as the activation function. MaxPooling1D layer with pool size $=2$ is added after CNN1D layer.
  • Transformer: The standard Transformer encoder Vaswani et al. [2017] is used. In our case, we only used 1 Encoder with 1 head. We set head size=128, embedding dimension=64, Dense layer with 32 neurons. Dropout is used after each layer above.
  • GTN: We combine the components of GNNs, LSTMs, Transformer, CNNs, and attention, as described in Fig. 5. We have $32$ and $16$ for GNN’s channels, $16$ for LSTM units, $96$ for CNN filters, and one Transformer encoder with $3$ heads (head size=192, embedding dimension=96).
Note.
When we used the above models for Flood Evaluator and Flood Manager, the input and output need to be changed accordingly (see Figures 2 and 3). To align with real-world scenarios, we enforce stringent constraints that restrict the output of Manager (values for gate and pump openings) to fall within the range of $[MIN,MAX]$ by a modified RELU activation function in Eq. (7). Fig. 10 provides the corresponding visualization.
$x=min\{max\{x,MIN\},MAX\},$ (7)
where $x$ is the opening value of gates and pumps.
Figure 10: Constraint for the values gate and pump openings.
B.2 Training hyper-parameters
In the following table, we show some hyper-parameters used in our experiments: learning rate, batch size, epoch, decay rate, decay step, patience for early stopping, dropout rate, and regularization factors $L_{1}$ and $L_{2}$.
Methods LR Batch Size Epoch Decay Rate Decay Step Patience Dropout Rate $L_{1}$ $L_{2}$
MLP 1e-3 512 3000 0.95 10,000 500 0.1 1e-5 1e-5
RNN 1e-4 512 3000 0.95 10,000 500 0.2 1e-5 1e-5
CNN 1e-4 512 3000 0.95 10,000 500 0.1 - -
GCN 5e-4 512 3000 0.95 10,000 500 0.0 - -
TCN 5e-4 512 3000 0.95 10,000 500 0.1 - -
RCNN 1e-4 512 3000 0.95 10,000 500 0.2 0 1e-5
Transformer 5e-4 512 3000 0.95 10,000 500 0.3 - -
GTN (ours) 5e-4 512 3000 0.95 10,000 500 0.5 1e-5 1e-5
Table 6: Best hyperparameters of each DL model for Flood Evaluator.
Methods LR Batch Size Epoch Decay Rate Decay Step Dropout Rate Patience $L_{1}$ $L_{2}$
MLP 1e-3 512 700 0.95 10,000 0.0 100 1e-5 1e-5
RNN 1e-3 512 700 0.95 10,000 0.2 100 1e-5 1e-5
CNN 3e-3 512 700 0.95 10,000 0.0 100 0.0 0.0
GNN 5e-4 512 700 0.9 10,000 0.0 100 0.0 0.0
TCN 1e-3 512 700 0.9 10,000 0.0 100 0.0 0.0
RCNN 1e-3 512 700 0.9 10,000 0.0 100 0.0 1e-5
Transformer 5e-4 512 700 0.9 10,000 0.5 100 0.0 0.0
GTN (ours) 3e-3 512 700 0.95 10,000 0.5 100 1e-5 1e-5
Table 7: Best hyperparameters of each DL model for Flood Manager.
Appendix C More experimental results
C.1 Flood prediction at all locations
In Table 2, we provided the comparison of the performance of different models for the Flood Evaluator at time point $t+1$, but at one location (S1). In Table 8, we provide the comparison of the performance of different models for the Flood Evaluator on the test set at time $t+1$, but for all locations.
Methods MAE (ft) RMSE (ft) Over Timesteps Over Area Under Timesteps Under Area
Ground-truth - - 427 68.61 5,116 1465.10
HEC-RAS 0.185 0.238 334 50.85 4,469 1274.26
MLP 0.060 0.080 900 245.75 5,493 1536.22
RNN 0.053 0.071 377 57.32 4,813 1308.28
CNN 0.075 0.105 342 53.69 5,182 1394.05
GCN 0.052 0.069 493 81.28 5,170 1454.78
TCN 0.063 0.097 335 48.37 5,479 1470.41
RCNN 0.121 0.138 288 45.10 5,631 1555.63
Transformer 0.049 0.065 515 85.49 5,046 1372.50
GTN (ours) 0.047 0.064 439 76.01 5,859 1689.45
Table 8: Comparison of different models for the Flood Evaluator on the test set (at time t+1 for all stations). “Over Timesteps” (and “Under Timesteps”) represent the number of time steps during which the water levels exceed the upper threshold (subceed the lower threshold, resp.). Similarly, “Over Area” (and “Under Area”) refer to the area between the water level curve and the upper threshold (lower threshold, resp.), as shown in Fig. 4 in the manuscript.
C.2 Flood mitigation at all locations
In Table 3, we provided the comparison of the performance of different models for the Flood Manager at time point $t+1$, but at one location (S1). We provide the comparison results of different models for the Flood Manager on the test set (at time t+1 for all locations) in Table 9.
Method Manager Over Timesteps Over Area Under Timesteps Under Area
Rule-based - 427 68.61 5,116 1465.10
GA-based Genetic $*{}^{*}$ (GA) - - - -
Genetic ${}^{{\dagger}}$ (GA) 343 67.50 1,817 416
DL-based MLP 435 67.28 3,511 818.06
RNN 211 25.28 386 58.39
CNN 372 54.72 3,969 1,044.71
GCN 181 25.06 1,109 207.48
TCN 224 25.61 562 92.16
RCNN 170 20.01 542 88.72
Transformer 374 54.31 4,378 1,142.98
${}^{{\ddagger}}$ (Ours) 114 12.09 617 98.42
Table 9: Comparison of different methods for the Flood Manager on the test set (at time t+1 for all stations). The $*{}^{*}$ denotes that the GA method was used for Flood Manager with a physics-based (HEC-RAS) as an Evaluator. -- denotes too long to get results. The ${}^{{\dagger}}$ denotes that the GA method was used for Flood Manager with the DL-based GTN as the Evaluator. All other rows are DL-based flood managers with a DL-based GTN as the Evaluator. ${}^{{\ddagger}}$ denotes the best.
C.3 Flood mitigation with one event
In Fig. 7, we visualize the results of different models for the Flood Manager on a short event spanning 18 hours from September 3rd (09:00) to September 4th (03:00) in 2019 for one location of interest. Below are the corresponding experiment results.
Method Manager Over Timesteps Over Area Under Timesteps Under Area
Rule-based - 6 0.866 0 0
GA-Based Genetic $*{}^{*}$ (GA) 4 0.351 0 0
Genetic ${}^{{\dagger}}$ (GA) 6 0.764 0 0
DL-Based MLP 6 0.614 0 0
RNN 1 0.074 0 0
CNN 6 0.592 0 0
GNN 2 0.062 0 0
TCN 1 0.046 0 0
RCNN 1 0.045 0 0
Transformer 6 0.614 0 0
${}^{{\ddagger}}$ (Ours) 1 0.022 0 0
Table 10: Comparison of different methods of Flood Manager for flood Mitigation (at time t+1 for one event S1).
Appendix D Visualization
Fig. 11 presents the results for all locations compared to Fig. 7 for one location of interest. Table 11 shows the corresponding performance measures.
Figure 11: Visualization for flood mitigation at all locations. Two gray dashed lines represent the upper (3.5 ft) and lower threshold (0.0 ft).
Method Manager Over Timesteps Over Area Under Timesteps Under Area
Rule-based - 23 3.62 0 0
GA-Based Genetic $*{}^{*}$ (GA) 16 1.39 0 0
Genetic ${}^{{\dagger}}$ (GA) 23 3.75 0 0
DL-Based MLP 23 4.18 0 0
RNN 8 0.63 0 0
CNN 21 2.92 0 0
GCN 6 0.32 0 0
TCN 6 0.39 0 0
RCNN 6 0.39 0 0
Transformer 22 3.00 0 0
${}^{{\ddagger}}$ (Ours) 5 0.22 0 0
Table 11: Comparison of different methods of Flood Manager for flood Mitigation (at time t+1 for one event at all locations).

### v2

@@
 Abstract
-In coastal river systems, frequent floods, often occurring during major storms or king tides, pose a severe threat to lives and property.
-However, these floods can be mitigated or even prevented by strategically releasing water before extreme weather events with hydraulic structures such as dams, gates, pumps, and reservoirs.
-A standard approach used by local water management agencies is the “rule-based” method, which specifies predetermined pre-releases of water based on historical and time-tested human experience, but which tends to result in excess or inadequate water release.
-The model predictive control (MPC), a physics-based model for prediction, is an alternative approach, albeit involving computationally intensive calculations.
-In this paper, we propose a Forecast Informed Deep Learning Architecture, FIDLAr, to achieve rapid and optimal flood management with precise water pre-releases.
-FIDLAr seamlessly integrates two neural network modules: one called the Flood Manager, which is responsible for generating water pre-release schedules, and another called the Flood Evaluator, which assesses these generated schedules.
-The Evaluator module is pre-trained separately, and its gradient-based feedback is used to train the Manager model, ensuring optimal water pre-releases.
-We have conducted experiments using FIDLAr with data from a flood-prone coastal area in South Florida, particularly susceptible to frequent storms.
+In coastal river systems, floods, often during major storms or king tides, severely threaten lives and property.
+However, hydraulic structures such as dams, gates, pumps, and reservoirs exist in these river systems, and these floods can be mitigated or even prevented by strategically releasing water before extreme weather events.
+A standard approach used by local water management agencies is the “rule-based” method, which specifies predetermined water prereleases based on historical human experience, but which tends to result in excessive or inadequate water release.
+Iterative optimization methods that rely on detailed physics-based models for prediction are an alternative approach.
+Whereas, such methods tend to be computationally intensive, requiring hours or even days to solve the problem optimally.
+In this paper, we propose a Forecast Informed Deep Learning Architecture, FIDLAr, to achieve rapid and near-optimal flood management with precise water prereleases.
+FIDLAr seamlessly integrates two neural network modules: one called the Flood Manager, which is responsible for generating water pre-release schedules, and another called the Flood Evaluator, which evaluates those generated schedules.
+The Evaluator module is pre-trained separately, and its gradient-based feedback is utilized to train the Manager model, ensuring near-optimal water pre-releases.
+We have conducted experiments with a flood-prone coastal area in South Florida.
 Results show that FIDLAr is several orders of magnitude faster than currently used physics-based approaches while outperforming baseline methods with improved water pre-release schedules.
-Our code is at https://github.com/JimengShi/FIDLAR/.
+Codehttps://github.com/JimengShi/FIDLAR
 1 Introduction
-Floods can result in catastrophic consequences with considerable loss of life Jonkman and Vrijling (2008), huge socio-economic impact Wu et al.
-(2021), property damage Brody et al.
-(2007), and environmental devastation Yin et al.
-(2023a).
-They pose a threat to food and water security as well as sustainable development Kabir and Hossen (2019).
-What is even more alarming is that research indicates global climate change may lead to a drastic increase in flood risks in terms of both frequency and scale Wing et al.
-(2022); Hirabayashi et al.
-(2013).
-In coastal areas, flood risks are increasing due to sea-level rise Sadler et al.
-(2020).
-Therefore, effective flood management is of utmost importance.
-To mitigate flood risks, water management agencies have built controllable hydraulic structures such as dams, gates, pumps, and reservoirs in river systems Kerkez et al.
-(2016).
-Determining optimal control schedules of these hydraulic structures is however a challenging problem Bowes et al.
-(2021).
-Rule-based methods Bowes et al.
-(2021); Sadler et al.
-(2019) formulate control schedules based on insights gained from historically observed data.
-These schedules are often formalized and extensively documented in databases Marchese et al.
-(2018).
-They represent the collective wisdom gathered over decades of experience in managing specific river systems and are commonly adopted by local water management authorities.
-However, these schedules may expose vulnerabilities during extreme events, which are becoming more likely due to climate change, and might lack sufficient locally available historical data, particularly for events like 500- or 1000-year floods.
-Moreover, these schedules may not offer effective solutions for complex river systems Schwanenberg et al.
-(2015).
-Model Predictive Control (MPC) is a well-known strategy that tackles this challenge as an optimization process Karimanzira (2016).
-By optimizing variables such as water levels and discharges, the algorithm generates control schedules for hydraulic structures to effectively mitigate overall flood risks Delaney et al.
-(2020); Zarei et al.
-(2021).
-Random initialization followed by soft-computing techniques such as genetic algorithms is often employed to perform the optimizations.
-Subsequently, physics-based models such as HEC-RAS and SWMM are used to assess the generated control schedules.
-Publications on this topic include the following Leon et al.
-(2020); Sadler et al.
-(2019); Vermuyten et al.
-(2018); Chen et al.
-(2016); Leon et al.
-(2014); Yin et al.
-(2023b).
-However, such MPC methods are prohibitively slow since they require thousands of simulations using physics-based models.
-In the last decade or so, Machine learning (ML) has emerged as a powerful tool in this domain.
-So far, the ML-based methods have been applied to flood prediction Mosavi et al.
-(2018); Shi et al.
-(2023b), flood detection Tanim et al.
-(2022); Shahabi et al.
-(2020), susceptibility assessment Saha et al.
-(2021); Islam et al.
-(2021), and post-flood management Munawar et al.
-(2019).
-However, ML-based methods have not been applied to flood mitigation tasks.
-We aim to fill this gap.
-Moreover, we replace the compute-intensive MPC approach with a framework based purely on deep learning (DL), where both the generator and evaluator are achieved by DL modules.
-Gradient-based planning with world models aims to seek optimal actions by training a world model and conducting inferences using that trained model.
-We refer the reader to Figure 1(a) from Jyothir et al.
-(2023).
+Floods can result in catastrophic consequences with considerable loss of life (Jonkman and Vrijling 2008), huge socio-economic impact (Wu et al.
+2021), property damage (Brody et al.
+2007), and environmental devastation (Yin et al.
+2023a).
+They pose a threat to food and water security as well as sustainable development (Kabir and Hossen 2019).
+What is even more alarming is that research indicates global climate change may lead to a drastic increase in flood risks in terms of both frequency and scale (Wing et al.
+2022), e.g., coastal flood risks due to sea-level rise (Sadler et al.
+2020).
+Thus, effective flood management is of utmost importance.
+To mitigate flood risks, water management agencies have built controllable hydraulic structures such as dams, gates, pumps, and reservoirs in river systems (Kerkez et al.
+2016).
+However, determining the optimal control schedules of these hydraulic structures is a challenging problem (Bowes et al.
+2021).
+Rule-based methods (Sadler et al.
+2019) formulate control schedules based on insights gained from historically observed data.
+The rules represent the collective wisdom gathered over decades of experience in managing specific river systems.
+Nevertheless, these rules may expose vulnerabilities while dealing with extremely rare events and may not offer effective solutions for complex river systems under such conditions (Schwanenberg, Becker, and Xu 2015).
 Figure 1: Forecast-Informed Deep Learning Architecture (FIDLAr).
-Input data in the form of a table consisting of five variables as shown in the left end of the diagram.
+Input data consists of five categories of variables as shown in the left table.
 The variables $w$ and $k$ are the lengths of the past and prediction windows, respectively.
-The parts of the table colored green are provided as inputs to the programs, while the orange and blue parts (with question marks) are outputs.
-The black and red arrows represent the propagation and back-propagation processes, respectively.
-The Flood Manager and Flood Evaluator represent DL models, the former to predict the control schedules of gates and pumps to pre-release water, and the latter to determine water levels for those control schedules.
+The parts colored green are provided as inputs, while the orange and blue parts (with question marks) are outputs.
+The Flood Manager and Flood Evaluator represent deep learning (DL) models, the former to predict control schedules of controllable hydraulic structures (e.g., gates and pumps) to pre-release water, and the latter to predict the resulting water levels for those control schedules.
 Loss functions, $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, penalize the flooding and water wastage beyond pre-specified thresholds, respectively.
-Our proposed Forecast Informed Deep Learning Architecture, FIDLAr, uses gradient-based planning to achieve optimal flood management with rapid and feasible water pre-releasing.
-See Figure 1 for details on its architecture.
+The flood mitigation problem is usually treated as an optimization task (Karimanzira 2016), where variables such as control schedules for hydraulic structures are optimized to achieve the desired water levels to effectively mitigate flood risks (Zarei et al.
+2021).
+Random initialization followed by soft-computing techniques such as genetic algorithm (GA) (Leon et al.
+2020) is often employed to perform the optimizations.
+Subsequently, physics-based simulator models such as HEC-RAS and SWMM are used to assess generated control schedules (Sadler et al.
+2019; Leon et al.
+2014; Yin et al.
+2023b).
+However, such methods are prohibitively slow since they require thousands of time-consuming physics-based simulations (Jafarzadegan et al.
+2023).
+Over the past decade, machine learning (ML) has become an increasingly powerful tool in various aspects of flood management, including flood prediction (Shi et al.
+2023c), flood detection (Tanim et al.
+2022), susceptibility assessment (Saha et al.
+2021), and post-flood management (Munawar et al.
+2019).
+Despite these advances, ML-based methods have yet to be widely applied to flood mitigation tasks.
+We address this gap by introducing a fully machine learning-based framework, comprising a generator to generate control schedules and a simulator to assess them.
+The ML-based framework offers significantly faster response times than compute-intensive physics-based approaches, enabling real-time flood management.
+Furthermore, the proposed ML-based method allows the simulator to actively guide the generator through gradient descent, thus seamlessly integrating the generator and simulator components and leveraging the differentiability of the learned simulator model (Jyothir et al.
+2023).
+To this end, we propose Forecast Informed Deep Learning Architecture, FIDLAr, for flood management.
 FIDLAr’s characteristics are summarized as follows:
@@
 The former model is responsible for generating water pre-release schedules, while the latter model accurately forecasts the resulting water levels.
-The latter model serves as an evaluator for the output of the former.
-During training, the gradient feedback from the Evaluator model can be back-propagated to reinforce the Manager model to generate effective water pre-release schedules.
-The Manager and Evaluator play the role of Planner and World Model, respectively, from the model from Jyothir et al.
-(2023).
+Moreover, with the gradient-based planning, and differentiability of trained Evaluator, it can reinforce the Manager to generate better schedules.
   • FIDLAr is a data-driven approach, learning flood mitigation strategies from historically observed data.
-Once trained, it offers rapid response capabilities.
-This highlights the advantages of DL-based models over physics-based models, particularly in the context of real-time flood management.
-  • FIDLAr is a model-agnostic framework, where both the Manager and Evaluator could be any type of DL model with differentiable loss functions that allow back-propagation.
-  • FIDLAr is trained with a custom loss function to balance both flood risks and water wastage at the same time, which is a multitask learning method.
-  • FIDLAr makes effective use of reliable forecasts of specific variables (e.g., precipitation and tidal information) for the near future.
-  • The explainability features of FIDLAr allow us to better understand the behavior of complex river systems.
+Once trained, it offers rapid response capabilities, highlighting the advantages of DL-based models over physics-based models, particularly for real-time flood management.
+  • FIDLAr is a model-agnostic framework, where both the Manager and Evaluator could be any type of DL model trained with differentiable loss functions that allow back-propagation.
+  • FIDLAr is trained with a customized loss function to balance flood risks and water wastage at the same time, which is a multitask learning method.
+2 Related Work
+Flood Prediction.
+Physics-based mechanic models (e.g., HEC-RAS, SWMM) have been widely used to simulate water levels and flows in river systems (Peker et al.
+2024; Rahman and Ali 2024; Gomes Jr et al.
+2023; Rivett et al.
+2022).
+However, these models are computationally inefficient and fall short of capturing precise knowledge of study domains (Bentivoglio et al.
+2022).
+Therefore, diverse machine learning (ML) and deep learning (DL) models have been studied as surrogates to simulate water levels and flows.
+For example, support vector machines (SVMs) have been used to predict the urban flash floods (Yan et al.
+2018; Choubin et al.
+2019), multivariate regression models were adopted to estimate flood volumes and peak flows (Yang and Chang 2020), random forests and K-nearest neighbors were explored for urban flood inundation mapping (Castro-Gama et al.
+2014), gaussian process learning models for fast and accurate flood inundation simulation (Fraehr et al.
+2023).
+Furthermore, deep learning models, such as recurrent neural networks, convolutional neural networks, and Transformers, have been employed for flood inundation (Zhou et al.
+2021) and flood prediction (Shi et al.
+2023b, c).
+Flood Mitigation.
+Flood mitigation seeks control schedules of those hydraulic structures to avoid or mitigate flood risks, which requires flood prediction models as simulators to evaluate the control schedules.
+Researchers have attempted to leverage the genetic algorithm and the pattern search to generate control schedules and physics-based models (e.g., EPA-SWMM5, HEC-RAS) as water simulators (Sadler et al.
+2019; Leon et al.
+2014, 2020).
+The Lake Mendocino Operations (LMO) model was developed to simulate operations of Lake Mendocino such as release constraints for flood control and water supply operations (Delaney et al.
+2020).
+However, such methods are computationally intensive since they require thousands of time-consuming simulation trials with physics-based models (Jafarzadegan et al.
+2023).
+Additionally, the genetic algorithm and pattern search techniques are usually heuristic, without positive feedback or guidance to better control those hydraulic structures.
 3 Problem Formulation
-Flood mitigation involves the prediction of control schedules for hydraulic structures like gates and pumps within the river system, denoted as $X^{gate,pump}_{t+1:t+k}$, spanning $k$ time points from $t+1$ to $t+k$ into the future.
-This prediction uses as input historical input data (on all variables), $X$, from the preceding $w$ time points, in conjunction with reliably forecasted covariates (such as rainfall and tide) for the next $k$ time points.
-Thus, we train a DL model whose transfer function, $\mathcal{F}_{\theta}$, has the form:
-$\mathcal{F}_{\theta}:(X_{t-w+1:t},X^{cov}_{t+1:t+k})\rightarrow X^{gate,pump}_{t+1:t+k},$ (1)
+Flood management or mitigation aims to manage water levels before extreme weather events.
+It involves predicting control schedules for hydraulic structures such as gates and pumps within the river system, denoted as $X^{gate,pump}_{t+1:t+k}$, spanning $k$ time points into the future from $t+1$ to $t+k$.
+This prediction takes as input historical data on all possible variables (see Figure 5), $X$, from the preceding $w$ time points, in conjunction with reliably forecasted covariates (such as rainfall and tide) for the next $k$ time points.
+Then we could train a deep learning (DL) model, $\mathcal{M}_{\theta}$, with parameters $\theta$:
+$\mathcal{M}_{\theta_{M}}:(X^{all}_{t-w+1:t},X^{cov}_{t+1:t+k})\rightarrow X^{gate,pump}_{t+1:t+k},$ (1)
 where the subscripts represent the time ranges, and the superscripts refer to the variables under consideration.
@@
 4 Methodology
-3.1 The Challenge
-Intuitively, an ML model can be trained to learn the function $f_{\theta}$.
-However, the challenge here is that the historical data may not represent the desired values for the control schedules of the gates and pumps, especially if it resulted in flood conditions, and therefore cannot be used as the ground-truth values used in traditional supervised learning.
-Therefore, we separate a “planner” (Flood Manager) module from an “evaluator” (Flood Evaluator) module.
-We first build an independent and accurate evaluator with the historical data, which attempts to learn the consequences of every possible action.
-We then treat the planner module design as an optimization problem where we optimize a custom loss function, allowing the system to pick the action that would prevent undesirable consequences.
-Desired states of the river systems avoid floods on one hand while preventing water wastage on the other.
-Once the Flood Evaluator is trained, it is frozen and used to train the Flood Manager.
-The evaluator serves to backpropagate the gradient feedback regarding the quality of the schedules generated by the manager, thus guiding the manager to generate more effective schedules of gates and pumps.
-Such a training process ends up refining the function, $f_{\theta}$, as represented in Eq.
-(1).
-Figure 2: Flood Evaluator.
-Unlike Figure 1, the parts shaded green (includes historical data and covariates predicted from near future) and orange (representing any control schedule for the gates and pumps) are both used as inputs.
-Water levels (blue) are the outputs.
+4.1 Overview
+Intuitively, an ML model can be trained to learn the function $\mathcal{M}_{\theta_{M}}$.
+However, a key challenge lies in the historical data, which often reflects control schedules that led to flooding or other suboptimal outcomes, making it unsuitable as ground-truth data for traditional supervised learning.
+To overcome this, we train the model, $\mathcal{M}_{\theta_{M}}$, by leveraging the differentiability of a learned simulator model, $\mathcal{E}_{\theta_{E}}$.
+First, we pre-train an independent and accurate Flood Evaluator using extensive historical data to model the consequences (e.g., water levels) of various actions (e.g., control schedules).
+We then frame the control schedule planning in Flood Manager as an optimization problem, seeking actions that minimize undesirable outcomes - floods or water wastage.
+Both the Evaluator and Manager are implemented as neural networks, with the framework illustrated in Figure 1.
 4.2 Flood Evaluator
-Flood Evaluator is tasked with accurately forecasting water levels at designated points of interest within river systems for any control schedule of gates and pumps provided to it.
-See Figure 2.
-The underlying transfer function of the Evaluator has the form:
-$\mathcal{G}_{\theta}:(X_{t-w+1:t},X^{cov}_{t+1:t+k},X^{gate,pump}_{t+1:t+k})\rightarrow X^{water}_{t+1:t+k}.$ (2)
-As mentioned earlier, the Evaluator module is trained independently using the historical data to achieve highly accurate water level predictions for any given set of conditions and control schedules.
-Therefore, once the Evaluator is trained, its parameters are set in stone for the training of the Manager, where it plays the role of a trained “referee”.
-Figure 3: Flood Manager. As in Figure 1, the parts shaded green (historical data) are the inputs, and the parts shaded orange are the outputs. The water levels shaded blue are not predicted.
+Flood Evaluator is tasked with accurately forecasting water levels at designated points of interest within river systems for any control schedule of gates and pumps.
+The underlying transfer function of the Evaluator is:
+$\mathcal{E}_{\theta_{E}}:(X^{all}_{t-w+1:t},X^{cov}_{t+1:t+k},X^{gate,pump}_{t+1:t+k})\rightarrow X^{water}_{t+1:t+k}.$ (2)
+The Evaluator is trained independently using large-scale historical data to achieve highly accurate water level predictions for any given set of conditions and control schedules.
+Therefore, once the Evaluator is well trained, its parameters are frozen while training the Manager, where it plays the role of a trained “referee” - scoring control schedules generated by the Manager by predicting the resulting water levels.
+It also serves to backpropagate the gradient descent feedback, guiding the Manager to produce more effective control schedules of gates and pumps.
+Figure 2: Flood Evaluator. The parts shaded green are used as inputs. Water levels (blue) are the outputs.
+Algorithm 1 Training algorithm of FIDLAr
+Input: recent past data: ${\bf X}^{all}_{t-w+1,t}$
+Input: near future data: ${\bf X}^{cov}_{t+1,t+k}={\bf X}^{rain,tide,gate,pump}_{t+1,t+k}$
+Parameter: $\theta_{E},\theta_{M}$: parameters of Evaluator and Manager; $w,k:$ length of past and prediction windows
+1:  // Train Flood Evaluator, $\mathcal{E}_{\theta_{E}}$
+2:  initialize learnable parameters $\theta_{E}$
+3:  for $i=1,\ldots,N$ epochs do
+4:     $\text{MiniBatch}\leftarrow(\{{\bf X}^{all}_{t-w+1,t},{\bf X}^{cov}_{t+1,t+k}\},{\bf X}^{water}_{t+1,t+k})$
+5:     $\hat{{\bf X}}^{water}_{t+1,t+k}\leftarrow\mathcal{E}_{\theta_{E}}({\bf X}^{all}_{t-w+1,t},{\bf X}^{cov}_{t+1,t+k})$
+6:     $\mathcal{L}_{E}\leftarrow\frac{1}{k}\sum_{j=1}^{k}||\hat{{\bf X}}^{water}_{j}-{\bf X}^{water}_{j}||^{2}$
+7:     $\nabla_{\theta_{E}}\leftarrow\text{BackwardAD}(\mathcal{L}_{E})$
+8:     $\theta_{E}\leftarrow\theta_{E}-\eta\nabla_{\theta_{E}}$
+9:  end for
+10:  return trained Flood Evaluator, $\mathcal{E}_{\theta_{E}}$
+11:  // Train Flood Manager, $\mathcal{M}_{\theta_{M}}$, with frozen $\mathcal{E}_{\theta_{E}}$
+12:  initialize learnable parameters $\theta_{M}$
+13:  while $X^{water}_{t,t+k}$ violates either threshold do
+14:     $\text{MiniBatch}\leftarrow(\{{\bf X}^{all}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k}\},{\bf X}^{gate,pump}_{t+1,t+k})$
+15:     $\hat{{\bf X}}^{gate,pump}_{t+1,t+k}\leftarrow\mathcal{M}_{\theta_{M}}({\bf X}^{all}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k})$
+16:     $\hat{{\bf X}}^{water}_{t+1,t+k}\leftarrow\mathcal{E}_{\theta_{E}}({\bf X}^{all}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k},\hat{{\bf X}}^{gate,pump}_{t+1,t+k})$
+17:     $\mathcal{L}_{E}=c_{1}\cdot\mathcal{L}_{1}(\hat{{\bf X}}^{water}_{t+1,t+k})+c_{2}\cdot\mathcal{L}_{2}(\hat{{\bf X}}^{water}_{t+1,t+k})$
+18:     $\nabla_{\theta_{M}}\leftarrow\text{BackwardAD}(\mathcal{L}_{E})$
+19:     $\theta_{M}\leftarrow\theta_{M}-\eta\nabla_{\theta_{M}}$
+20:  end while
+21:  return trained Flood Manager, $\mathcal{M}_{\theta_{M}}$
 4.3 Flood Manager
-As shown in Figure 3, the Flood Manager is trained to produce optimal control schedules for the hydraulic structures (gates and pumps), taking reliably predictable future information (rain, tide) and all historical data.
-During its training, since no ground truth is available, it is trained using the gradient descent algorithm Ruder (2016); Andrychowicz et al.
-(2016), using the partial derivatives of the loss function (Eq.
-(3)), which in turn needs the evaluator.
+Flood Manager is to produce control schedules for hydraulic structures (i.e., gates and pumps), taking as inputs reliably predictable future information (rain, tide) and all historical data.
+Since no ground truth is available, it is trained with the differentiability of the learned Evaluator model.
+Therefore, we connect the Manager with the Evaluator where the output of Eq.
+(1) is injected into Eq.
+(2):
+$\mathcal{E}_{\theta_{E}}(X^{all},X^{cov}_{t+1:t+k},\mathcal{M}_{\theta_{M}}(X^{all},X^{cov}_{t+1:t+k}))\rightarrow X^{water}_{t+1:t+k},$ (3)
+where $X^{all}=X^{all}_{t-w+1:t}$ and $\theta_{M}$ and $\theta_{E}$ are the parameters of Manager and Evaluator.
+The resulting output of water levels can be used to compute the loss in Eq.
+(6), representing evaluation scores for generated control schedules.
+Gradient descent (Ruder 2016) can be back-propagated as the feedback to update the parameters of the Manager.
+The parameter update is presented:
+$\theta_{M}:=\theta_{M}-\alpha\cdot\frac{\partial\mathcal{L}}{\partial\theta_{M}},$ (4)
+where $\alpha$ is the learning rate and $\frac{\partial\mathcal{L}}{\partial\theta_{M}}$ is the partial derivative of the compound function $\frac{\partial\mathcal{L}}{\partial\theta_{M}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}}\cdot\frac{\partial\mathcal{E}}{\partial\mathcal{M}}\cdot\frac{\partial\mathcal{M}}{\partial\theta_{M}}$.
+The training details of FIDLAr are given in Algorithm 1.
+Figure 3: Flood Manager. The parts shaded green (historical data) are the inputs, and the parts shaded orange are the outputs. The water levels shaded blue are not predicted.
 4.4 Custom Loss Function
 Loss functions are critical in steering the learning process.
-For flood management, since no ground truth is available, independent metrics of performance are needed.
 An obvious one is the total time (Figure 4(a)) for which the water levels either exceed the flooding threshold or dip below the water wastage threshold.
@@
 It also prevents the optimization methods from trivially recommending the depletion of valuable water resources to prevent future flooding.
-The above metrics suggest the following components for the loss function for our models.
-The final loss function is a balanced combination of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ in Eq.
+The $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ represent the flooding and water wastage losses, respectively, and the final loss function is a balanced combination as shown in Eq.
 (6).
@@
 $\mathcal{L}_{2}=\sum^{N}_{i=1}\sum^{t+k}_{j=t+1}\Arrowvert min\{\hat{X}^{water}_{i,j}-X^{waste}_{i},0\}\Arrowvert^{2},$
-where $N$ is the number of water level locations of interest; $k$ is the length of prediction horizon; $X^{flood}$ and $X^{waste}$ represent the thresholds for flooding and water wastage; and their capped versions, $\hat{X}^{flood}$ and $\hat{X}^{waste}$ are obtained using the evaluator module.
+where $N$ is the number of water level locations of interest; $k$ is the length of prediction horizon; $X^{flood}$ and $X^{waste}$ represent the thresholds for flooding and water wastage; and the capped version, $\hat{X}^{water}$, is obtained using the Evaluator module.
 The combined loss function is given by:
-$\mathcal{L}_{total}=\alpha\cdot\mathcal{L}_{1}+\beta\cdot\mathcal{L}_{2},$ (4)
-where $\alpha/\beta$ dictates the relative importance of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$.
-Figure 4: The two red bars represent an upper threshold of flooding and a lower threshold of wastage.
-Shown are (a) the time spans when these thresholds are crossed, and (b) the areas between the water level curve and the threshold bars.
-Violations of the upper (lower, resp.) threshold are captured in $\mathcal{L}_{1}$ ($\mathcal{L}_{2}$, resp.).
-Figure 5: Architecture of Graph Transformer Network for Flood Evaluator.
-Input variables include Rainfall, Pump, Gate, Tide, and Water levels as shown in Figure 6 and Table 1, which are generically denoted by $R,T,G,P$, and $W$.
-Output is water levels.
-We use the same architecture for Evaluator and Manager, but with minor changes accordingly of the input and output(see Figs.
-2 and 3).
-3.5 Training FIDLAr
-The entire training process is outlined below.
-The Evaluator model is trained in a traditional way using the observed historical data as the ground truth for predictions and minimizing a loss function that penalizes deviations from the ground truth.
-After the Evaluator model is trained, its parameters are frozen and the Manager model is trained.
-For every control schedule that the Manager generates, $X^{gate,pump}_{t+1:t+k}$, the Evaluator is used to compute the resulting water levels, which is then used to compute the loss function defined in Eq.
-(4).
-Gradient descent Ruder (2016) provides the information to be back-propagated to be used to update the parameters of Manager.
-Thus,
-${X^{*}}^{gate,pump}_{t+1:t+k}=\underset{X^{gate,pump}_{t+1:t+k}}{\arg\min}\{\mathcal{L}_{total}\}.\$ (5)
-The training details of FIDLAr are given in Algorithm 1.
-4 GTN Model
+$\mathcal{L}_{total}=c_{1}\cdot\mathcal{L}_{1}+c_{2}\cdot\mathcal{L}_{2},$ (6)
+where $c_{1}/c_{2}$ dictates the relative importance of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$.
+Figure 4: The two red bars represent a threshold of flooding and a threshold of water wastage.
+Shown are (a) the time spans when these thresholds are crossed, and (b) the areas between water level curves and threshold bars.
+Violations of the upper and lower thresholds are captured in $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$.
+5 Graph Transformer Network
 The Manager and Evaluator modules described so far are model agnostic.
 We tried many existing architectures for them, as discussed in Section 6.2.
-The Graph Transformer Network (GTN) architecture combines graph neural networks (GNNs), attention-based transformer networks, long short-term memory networks (LSTMs), and the convolutional neural networks (CNNs).
-GNN and LSTM modules are combined for learning the spatiotemporal dynamics of water levels, while the Transformer and CNN modules focus on extracting feature representations from the covariates.
-The Attention mechanism is employed to discern the interactions between covariates and water levels, as shown in Eq.
+We devise the Graph Transformer Network (GTN) architecture by combining graph neural networks (GNNs), attention-based transformer networks, long short-term memory networks (LSTMs), and convolutional neural networks (CNNs).
+GNN and LSTM modules are combined to learn the spatiotemporal dynamics of water levels, while the Transformer and CNN modules focus on extracting feature representations from the covariates.
+The attention mechanism (Vaswani et al.
+2017) is used to discern interactions between covariates and water levels, as shown in Eq.
 (7) below.
-Figure 5 presents the GTN architecture used (with minor modifications) for the Evaluator and the Manager.
-$Atte(Q,K,V){}$ $=softmax(\frac{Q^{cov}(K^{water})^{T}}{\sqrt{d_{k}}})V^{water}$ (6)
-$=softmax(\frac{Q^{water}(K^{cov})^{T}}{\sqrt{d_{k}}})V^{cov},$
-where $T$ denotes the transpose operation; $water$ and $cov$ represent the variables of water levels and covariates.
-Algorithm 1 Training algorithm of FIDLAr
-Input: recent past data: ${\bf X}_{t-w+1,t}$
-Input: near future data: ${\bf X}^{cov}_{t+1,t+k}={\bf X}^{rain,tide,gate,pump}_{t+1,t+k}$
-Parameter: $\theta_{E},\theta_{M}$: parameters of Evaluator and Manager
-Parameter: $w,k:$ length of past and prediction windows
-1:  // Train Flood Evaluator, $\mathcal{G}_{\theta_{E}}$
-2:  initialize learnable parameters $\theta_{E}$
-3:  for $i=1,\ldots,N$ epochs do
-4:     $\text{MiniBatch}\leftarrow(\{{\bf X}_{t-w+1,t},{\bf X}^{cov}_{t+1,t+k}\},{\bf X}^{water}_{t+1,t+k})$
-5:     $\hat{{\bf X}}^{water}_{t+1,t+k}\leftarrow\mathcal{G}_{\theta_{E}}({\bf X}_{t-w+1,t},{\bf X}^{cov}_{t+1,t+k})$
-6:     $\mathcal{L}_{E}\leftarrow\frac{1}{k}\sum_{j=1}^{k}||\hat{{\bf X}}^{water}_{j}-{\bf X}^{water}_{j}||^{2}$
-7:     $\nabla_{\theta_{E}}\leftarrow\text{BackwardAD}(\mathcal{L}_{E})$
-8:     $\theta_{E}\leftarrow\theta_{E}-\eta\nabla_{\theta_{E}}$
-9:  end for
-10:  return trained Flood Evaluator, $\mathcal{G}_{\theta_{E}}$
-11:  // Train Flood Manager, $\mathcal{F}_{\theta_{M}}$, with frozen $\mathcal{G}_{\theta_{E}}$
-12:  initialize learnable parameters $\theta_{M}$
-13:  while $X^{water}_{t,t+k}$ violates either threshold do
-14:     $\text{MiniBatch}\leftarrow(\{{\bf X}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k}\},{\bf X}^{gate,pump}_{t+1,t+k})$
-15:     $\hat{{\bf X}}^{gate,pump}_{t+1,t+k}\leftarrow\mathcal{F}_{\theta_{M}}({\bf X}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k})$
-16:     $\hat{{\bf X}}^{water}_{t+1,t+k}\leftarrow\mathcal{G}_{\theta_{E}}({\bf X}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k},\hat{{\bf X}}^{gate,pump}_{t+1,t+k})$
-17:     $\mathcal{L}_{E}=\alpha\cdot\mathcal{L}_{1}(\hat{{\bf X}}^{water}_{t+1,t+k})+\beta\cdot\mathcal{L}_{2}(\hat{{\bf X}}^{water}_{t+1,t+k})$
-18:     $\nabla_{\theta_{M}}\leftarrow\text{BackwardAD}(\mathcal{L}_{E})$
-19:     $\theta_{M}\leftarrow\theta_{M}-\eta\nabla_{\theta_{M}}$
-20:  end while
-21:  return trained Flood Manager, $\mathcal{F}_{\theta_{M}}$
+Figure 6 presents the GTN architecture, which is used for both Evaluator and Manager, but with minor changes accordingly of the inputs and outputs (see Figures 2 and 3).
+We have two GNN layers with 32 and 16 channels, one LSTM layer with 16 units, one CNN with 96 filters, and one Transformer encoder with 3 heads.
+$Atte(Q,K,V){}$ $=softmax(\frac{Q^{cov}(K^{water})^{T}}{\sqrt{d}})V^{water}$ (7)
+$=softmax(\frac{Q^{water}(K^{cov})^{T}}{\sqrt{d}})V^{cov},$
+where $T$ denotes the transpose operation; $water$ and $cov$ represent water levels and covariates; and $d$ is the length of projection embedding where $d=d_{q}=d_{k}=d_{v}=128$.
 6 Experiments
-5.1 Data
-We obtained data from the South Florida Water Management District’s (SFWMD) DBHydro database District (2023) for the coastal stretch of the South Florida watershed.
-The data set consists of hourly observations for water levels and external covariates from Jan.
-1st, 2010 to Dec.
-31, 2020, a period totaling eleven years.
-As shown in Figure 6, the river system under consideration has two branches and includes several hydraulic structures (gates, pumps) to control water flows.
-Water levels in coastal systems are also impacted by tides.
-We aim to predict effective schedules on hydraulic structures (gates, pumps) to minimize flood risks at four specific locations marked by green circles in Figure 6.
-To highlight the risk posed by floods, we note that this portion of the river system flows through Miami downtown, with a sizable population, many commercial enterprises, and an international airport in its vicinity.
-Figure 6: Schemetic diagram of study domain - South Florida.
-There are three water stations with hydraulic structures (S26, S25B, S25A), one simple water station, S1 (green circle in the middle), and one station monitoring tide information from the ocean.
-Table 1: Summary of the data set (Jan. 1, 2010 $\sim$ Dec. 31, 2020).
+6.1 Dataset
+We obtained data from the South Florida Water Management District’s (SFWMD) DBHydro database (District 2023) for the coastal stretch in South Florida.
+The data set consists of hourly observations for water levels and external covariates from January 1, 2010 to December 31, 2020.
+As shown in Figure 5, the river system has two branches and includes several hydraulic structures (gates, pumps) to control water flows.
+We aim to predict effective control schedules on hydraulic structures (gates, pumps) to minimize flood risks at four specific locations marked by green circles.
+Figure 5: Schemetic diagram of study domain.
+There are three water stations with hydraulic structures (S26, S25B, S25A), one simple water station, S1 (green circle in the middle), and a station monitoring the tide level from the ocean.
 Feature Interval Unit #Var. Location
@@
 Water Hourly $ft$ 4 S25A, S25B, S26, S1
-5.2 Experimental designs
-The sliding input window Li et al.
-(2014) (also known as look-back window Gidea and Katz (2018) strategy was used to process the entire dataset Shi et al.
-(2023a)), which helps to ensure that the input and output vector lengths are constant.
-The datasets were split in chronological order with the first 80% for training and the remaining 20% for testing.
+Table 1: Summary of the data set.
+Figure 6: Graph Transformer Network for Flood Evaluator.
+Input variables include Rainfall, Pump, Gate, Tide, and Water levels as shown in Figure 5 and Table 1, which are generically denoted by $R,T,G,P$, and $W$.
+Output is water levels.
+6.2 Experimental Design
+The sliding input window (Li et al.
+2014) (also known as look-back window (Gidea and Katz 2018) strategy was used to process the entire dataset (Shi et al.
+2023a)).
+For consistency, we used a look-back window of length $w=72$ hours and a prediction window of length $k=24$ hours.
+The dataset was split in chronological order with the first 80% for training and the remaining 20% for testing.
+The eight DL methods below are used for Flood Manager and Flood Evaluator.
+We run all experiments on one NVIDIA A100 GPU with 80GB memory.
+  • MLP (Suykens, Vandewalle, and De Moor 1995): Multilayer perceptron can learn non-linear dependencies;
+  • RNN (Medsker and Jain 2001): Recurrent neural networks are good at processing sequential data;
+  • CNN (O’Shea and Nash 2015): A 1D convolutional neural network;
+  • GNN (Kipf and Welling 2016): Graph neural network with nodes representing variables and edges representing spatial dependencies;
+  • TCN (Bai, Kolter, and Koltun 2018): Temporal dilated convolutional network with an exponentially large receptive field;
+  • RCNN (Zhang and Dong 2020): Combined RNN and CNN model for time series forecasting;
+  • Transformer (Vaswani et al. 2017): Attention-based network for sequence modeling (only encoder);
+  • GTN (Ours): Combining GNNs with LSTMs, CNNs, and transformers, as described in Figure 6.
 Flood Prediction.
-The primary role of the Flood Evaluator is to forecast flood events by predicting water levels for given input conditions.
-To assess its performance, we compared several DL models for the evaluation, including our proposed GTN model as well as a physics-based model (HEC-RAS).
-We measured accuracy using multiple metrics: (a) the mean absolute error (MAE), (b) the root mean square error (RMSE) computed between the predicted and actual water levels, (c) the number of time points where the upper or lower thresholds were breached, and (d) the area between the water level curve and the threshold bars.
-For consistency, we used a look-back window of length $w=72$ hours and a prediction window of length $k=24$ hours.
-The eight DL methods used are listed below (more details in Appendix B):
-  • MLP Suykens et al. (1995): Multilayer perceptron can learn complex non-linear dependencies;
-  • RNN Medsker and Jain (2001): Recurrent neural networks are good at processing sequential data;
-  • CNN O’Shea and Nash (2015): A 1D convolutional neural network;
-  • GNN Kipf and Welling (2016): Graph neural network with nodes representing variables and edges representing spatial dependencies;
-  • TCN Bai et al. (2018): Temporal dilated convolutional network with an exponentially large receptive field;
-  • RCNN Zhang and Dong (2020): Combined RNN and CNN model for time series forecasting;
-  • Transformer Vaswani et al. (2017): Attention-based network for sequence modeling (only encoder);
-  • GTN: Combining GNNs with LSTMs, CNNs, and transformers, as described in Figure 5.
-Flood Mitigation with FIDLAr.
-FIDLAr requires both Evaluator and Manager components and we chose the GTN model for the Evaluator, since it was the best performing model.
-For the Manager model, we experimented with one rule-based method, and two genetic algorithms – one with a physics-based HEC-RAS evaluator Leon et al.
-(2020) and one with our DL-based GTN evaluator, and several DL-based managers.
-As shown in Table 3, the list of DL approaches used in our experiments for the Manager included MLP, RNN, CNN, GNN, TCN, RCNN, Transformer, and GTN.
-Performance of FIDLAr was measured using (a) the number of time steps where the upper/lower thresholds were exceeded for the water levels, and (b) the area between the water level curve and the threshold bars.
-5.3 Analysis of computational time
-Since FIDLAr was designed for real-time flood control, we measured the running times of the models used in this work.
-Table 2: Comparison of model performances for the Flood Evaluator on the test set, specifically at time t+1 for measurement station S1.
-The terms “Over Timesteps” and “Under Timesteps” indicate the number of time steps during which water levels exceed the upper threshold or fall below the lower threshold, respectively.
-Similarly, “Over Area” and “Under Area” pertain to the area between the water level curve and upper or lower threshold, as was illustrated in Figure 4.
-Results in red are lowest in that column or closest to the ground truth (in blue).
+The role of the Flood Evaluator is to forecast flood events by predicting water levels for given input conditions.
+We set the upper threshold (flood level) at 3.5 feet and the lower threshold (wastage level) at 0.0 feet.
+However, the methods remain consistent for many reasonable choices of threshold values.
+We measured accuracy using multiple metrics: (a) mean absolute error (MAE), (b) root mean squared error (RMSE) computed between the predicted and actual water levels, (c) number of time points where the upper or lower thresholds are breached, and (d) the area between water level curves and threshold bars.
+Table 2 demonstrates that our model GTN outperforms other models with predictions (in red) most closely aligned with the ground truth (in blue) while achieving the lowest MAE and RMSE (in orange).
+Therefore, we choose our GTN model as Evaluator while training Manager in FIDLAr.
+In the Appendix, we provide more results in Table 8.
 Methods MAE (ft) RMSE (ft) Over Timesteps Over Area Under Timesteps Under Area
@@
 GTN (ours) 0.040 0.056 100 15.64 1,390 398.84
-Table 3: Comparison of model performances for the Flood Manager on the test set, specifically at time t+1 for measurement station S1.
-The $*{}^{*}$ denotes that the GA method was used with a physics-based (HEC-RAS) evaluator.
-The -- denotes that the experiments were timed out.
-The ${}^{{\dagger}}$ denotes that the GA method was used with the GTN as the evaluator.
-All other rows are DL-based flood managers with a DL-based GTN as the evaluator.
-Results in red font are the best in that column.
+Table 2: Comparison of model performances for the Flood Evaluator on the test set, specifically at time t+1 for measurement station S1.
+The terms “Over Timesteps” and “Under Timesteps” indicate the number of time steps during which water levels exceed the upper threshold or fall below the lower threshold, respectively.
+Similarly, “Over Area” and “Under Area” pertain to the area between the water level curve and upper or lower threshold, as was illustrated in Figure 4.
+Results in orange are the lowest in that column while results in red are the closest to the ground truth (in blue).
+Flood Mitigation with FIDLAr.
+FIDLAr requires both Evaluator and Manager components.
+For the Manager model, we experimented with one rule-based method, and two genetic algorithms – one with a physics-based HEC-RAS evaluator (Leon et al.
+2020) and one with our DL-based GTN evaluator, and several DL-based managers using MLP, RNN, CNN, GNN, TCN, RCNN, Transformer, and GTN.
+FIDLAr was measured using (a) the number of time steps where the upper/lower thresholds are exceeded for the water levels, and (b) the area between the water level curves and the threshold bars.
+Table 3 shows that all DL-based methods consistently performed better for site S1 than rule-based and GA-based approaches.
+Furthermore, GTN has the best performance under all four metrics, whether it is to control floods or water wastage.
+The results for all other measurement sites follow a similar pattern, as shown in Table 9 in Appendix C.2.
 Method Manager Over Timesteps Over Area Under Timesteps Under Area
 Rule-based 96 14.82 1,346 385.8
-GA-based Genetic $*{}^{*}$ (GA) - - - -
-Genetic ${}^{{\dagger}}$ (GA) 86 16.54 454 104
+GA-based Genetic Algorithm - - - -
+Genetic Algorithm 86 16.54 454 104
 DL-based MLP 91 13.31 1,071 268.35
@@
 GTN (Ours) 22 2.23 299 53.34
-6 Experimental Results
-6.1 Flood Prediction
-To evaluate the accuracy of the Flood Evaluator, we test FIDLAr with various DL models on the test set by predicting the water levels $k=24$ hours in advance.
-Table 2 shows the MAEs and RMSEs for the DL-based models and the physics-based model (HEC-RAS).
-As discussed earlier, we also analyze the time steps and areas over and under the thresholds.
-We set the upper threshold (flood level) at 3.5 feet and the lower threshold (wastage level) at 0.0 feet.
-However, the methods remain consistent for many reasonable choices of threshold values.
-From Table 2, we note that GTN surpasses the performance of other models with predictions most closely aligned with the ground truth (highlighted in red).
-More results are provided in Table 8 in Appendix C.1.
-6.2 Flood Mitigation
-Performance measures.
-Table 3 shows that all DL-based methods consistently performed better for site S1 than rule-based and GA-based approaches.
-Moreover, GTN has the best performance under all four metrics, whether it is to control floods or water wastage.
-The results for all other measurement sites follow a similar pattern, as shown in Table 9 in Appendix C.2.
-Performance visualization.
-We visualize the mitigated water levels for a short sample spanning 18 hours from September 3rd (09:00) to September 4th (03:00) in 2019 for one location of interest.
-Figure 7 indicates that FIDLAr equipped with GTN has resulted in water levels within the upper and lower thresholds.
-The zoomed portion shows a 2.5-hour period where the water levels are decreased under 3.5 feet based on the predicted gate and pump schedules.
-We provide the corresponding performance in Table 10.
-More visuals are in Appendix D.
-Figure 7: Visualization of water levels with various methods for flood mitigation. We zoomed in the period of $t=2\sim 4.5$ in gray.
+Table 3: Comparison of model performances for the Flood Manager on the test set, specifically at time t+1 for measurement station S1.
+The denotes that the GA method was used with a physics-based (HEC-RAS) evaluator.
+The -- denotes that the experiments were timed out.
+The denotes the GA method was used with the GTN as the evaluator.
+All other rows are DL-based flood managers with a DL-based GTN as the evaluator.
+Results in bold are the best in that column.
+We visualize water levels for a short period spanning 18 hours from September 3rd (09:00) to September 4th (03:00) in 2019 for the S1 location.
+Figure 7 indicates that FIDLAr equipped with GTN model (purple curve) has led to water levels within the upper and lower thresholds, satisfying pre-defined requirements.
+Moreover, FIDLAr presents the best control (i.e., the least water levels beyond thresholds) for flood mitigation and water waste compared to other baselines.
+We zoomed in on a 2.5-hour period of the resulting decreased water levels.
+Its quantitative performance is in Table 10 while more visuals are in Appendix D.
+Figure 7: Visualization of water levels with various methods for flood mitigation. We zoomed in $t=2\sim 4.5$ in gray. ‘A’–‘B’ in legend represents the Manager and Evaluator.
 6.3 Ablation Study
-The ablation study quantifies the contribution of each of the components of GTN, by measuring the performance of GTN after removing each of the individual components. Results are shown in Table 4.
+The ablation study in Table 4 quantifies the contribution of each component of GTN, by measuring the performance of GTN after removing each of the individual components.
+6.4 Analysis of Computational Time
+Since FIDLAr was designed for real-time flood control, we measured the running times of the models used in this work.
+Table 5 shows the running times for the whole flood prediction and mitigation system in its training and test phase.
+All the DL-based approaches in the test phase are several orders of magnitude faster than the currently used physics-based and GA-based approaches for the flood mitigation task.
+Rapid inference is a critical property of data-driven DL methods.
+The table also shows the training times for the DL-based approaches, although they are not necessary for the deployment in reality.
+6.5 Explainability
+Attention-based methods allow us to calculate the “attention scores” assigned to an input variable to compute a specific output variable.
+This is exemplified in the heatmap shown in Figure 8, which shows the attention scores assigned to the tide (columns) to compute the gate schedule output (rows) for 24 hours into the future.
+Note that there are 96 columns and 24 rows because we use 72 hours of past tidal observations and 24 hours of future predicted tidal data to predict 24 hours of the gate schedule into the future.
+The rows $[0,23]$ correspond to the 24 hours into the future, while the columns $[0,95]$ also include 72 hours of the recent past and 24 hours of the future predicted tidal data.
+Therefore, $t=72$ corresponds to the “current” time point, and the columns $[72,95]$ correspond to the same time points as rows $[0,23]$.
+Figure 8: Importance scores of tide input. $x$ and $y$ axes are the tide and gate over time.
+Over Over Under Under
+Method Timesteps Area Timesteps Area
+w/o CNN 37 4.37 476 85.54
+w/o Transformer 32 3.57 325 57.42
+w/o GNN 56 5.90 479 86.22
+w/o LSTM 35 4.34 329 56.74
+w/o Attention 32 3.59 341 60.48
+GTN (Ours) 22 2.23 299 53.34
 Table 4: Ablation study for flood mitigation for the entire test set (for time point t+1 at S1).
 The last row indicates the performance of the FIDLAr system with GTN as proposed in Figure 6.
-The best results are in red font.
-Removed Over Over Under Under
-Method Timesteps Area Timesteps Area
-CNN 37 4.37 476 85.54
-Transformer 32 3.57 325 57.42
-GNN 56 5.90 479 86.22
-LSTM 35 4.34 329 56.74
-Attention 32 3.59 341 60.48
-GTN 22 2.23 299 53.34
-6.4 Computational time
-Table 5 shows the running times for all the methods for the evaluator component and for the whole flood mitigation system in its test phase.
-All the DL-based approaches are several orders of magnitude faster than the currently used physics-based and GA-based approaches.
-The table also shows the training times for the DL-based approaches, although they are not necessary for real-time performance.
-Table 5: The running time for flood prediction and mitigation.
-The running time for the rule-based method is not reported since historical data was directly used.
-$*{}^{*}$, which combines a GA-based tool and HEC-RAS, took too long and was not reported although the time for a small data set is reported in Tables 10, 11 in Appendix C.3.
-${}^{{\dagger}}$, which combines the GA-based tool with GTN, also took too long but was estimated using a smaller sample.
-Model Prediction Mitigation
+The best results in the last row are in bold.
+Model Flood Prediction Flood Mitigation
 Train Test Train Test
@@
 Rule-based - - - -
-${}^{{\dagger}}$ est. 30 h
+GA - - - -
+GA - - - est. 30 h
 MLP 35 min 1.88 s 58 min 6.13 s
@@
 Transformer 43 min 2.38 s 23 min 6.76 s
-GTN 119 min 2.95 s 35 min 4.90 s
-6.5 Model explainability
-Attention-based methods allow us to calculate the “attention scores” assigned to an input variable to compute a specific output variable.
-This is exemplified in the heatmap shown in Figure 8, which shows the attention scores assigned to the tide (columns) to compute the gate schedule output (rows) for 24 hours into the future.
-Note that there are 96 columns and 24 rows because we use 72 hours of past tidal observations and 24 hours of future predicted tidal data to predict 24 hours of the gate schedule into the future.
-Thus the rows $[0,23]$ correspond to the 24 hours into the future, while the columns $[0,95]$ also include 72 hours of the recent past.
-Thus, $t=72$ corresponds to the “current” time point, and the columns $[72,95]$ correspond to the same time points as rows $[0,23]$.
-Figure 8: Attention scores between tide input and one of the gate schedule outputs. $x$ and $y$ axes are the tide and gate over time.
+GTN (Ours) 119 min 2.95 s 35 min 4.90 s
+Table 5: Running time for flood prediction and mitigation.
+The running time for the rule-based method is not reported since historical data was directly used.
+GA, which combines a GA-based tool and HEC-RAS, took too long and was not reported, although the time for a small data set is reported in the appendix of our arXiv version.
+GA, which combines the GA-based tool with GTN, also took too long but was estimated using a smaller sample.
 7 Discussion
 Interpreting Attention Map.
-The explainability feature, which was shown with an example in Figure 8, can provide significant insights into our results.
-Firstly, the simple example with the attention paid to the variable “tide” shows that the brightest patches in the heatmap occur in the last 24 columns.
-This highlights the importance of the novel idea of using tidal forecasting information in our flood mitigation approach.
-Given that tidal information within a 24-hour window can be quite accurate, this is useful.
+The explainability feature, which is shown with an example in Figure 8, can provide significant insights into our results.
+Firstly, we point out that the brightest patches are in the last 24 columns of the heatmap.
+Thus FIDLAr pays greater attention to the 24 hours of future predicted tidal data than the past 72 hours, highlighting the importance of forecast information to a DL-based approach to flood mitigation.
 While tides may have a more predictable pattern over time, the contribution of rain on the water levels can also be seen for other time points.
@@
 Additionally, the water level at the first high tide peak after the “current” time is more significant than the other two.
-Third, the gate schedule has peaks at times $t=11h$ and $t=22h$ into the future, which correspond to the lowest points of the tide.
+Third, the gate schedule has peaks at times $t=11$ hour and $t=22$ hour into the future, which correspond to the lowest points of the tide.
 This implies that the optimal time for pre-releasing water is during low tide phases.
 Opening gates during high tide periods in coastal river systems is less advisable, as it may lead to water flowing back upstream from the ocean.
-Finally, we observe that while there is a light patch around column $t=65h$, suggesting mild attention for the previous high tide peak, but almost no attention to any of the peaks prior to that.
+Finally, we observe that there is a light patch around column $t=65$ hour, suggesting mild attention for the previous high tide peak, but almost no attention to any of the peaks before that.
 This again suggests that we could have chosen to use a smaller window for the past input.
-Doing this analysis for more time points could provide evidence for the right value of $w$, the size of the look-back window.
+Doing this analysis could provide evidence for the right value of $w$, the size of the look-back window.
 8 Conclusions
-FIDLAr is a DL-based tool to compute water “pre-release” schedules for hydraulic structures in a river system to achieve effective and efficient flood mitigation, while ensuring that water wastage is avoided.
+In this work, we present the shortcomings of the current approaches for flood mitigation and propose FIDLAr, a DL-based tool to address the problem.
+FIDLAr can compute water “pre-release” schedules for hydraulic structures in a river system to achieve effective and efficient flood mitigation, while ensuring that water wastage is avoided.
 This was made possible by the use of well-crafted loss functions for the DL models.
 The dual component design (with a Manager and an Evaluator) is a strength of FIDLAr.
-During training, the gradient-based back-propagation from the Evaluator helps to fine-tune the Manager.
-The reliability of forecasts of variables like precipitation and tides is well known, but their critical importance may have been best demonstrated for the first time using the attention mechanism employed herein.
+It exploits the gradient-based planning and the differentiability of the trained Evaluator model for better optimization.
+During training, the gradient-based back-propagation from the Evaluator helps to reinforce the Manager.
 All the DL-based versions of FIDLAr are several orders of magnitude faster than the (physics-based or GA-based) competitors while achieving improvement over other methods in flood mitigation.
-These characteristics allow us to entertain the possibility of real-time flood management, which was previously unthinkable.
+These characteristics allow us to entertain the possibility of real-time flood management, which was challenging for previous approaches.
+Broader Impact
+This research focuses on flood mitigation which has an essential influence on disaster management.
+As an AI application for social good, our model provides detailed responses on the management of hydraulic structures so as to avoid or mitigate flood events while ensuring minimal water wastage.
+Our model can also serve as a warning tool.
+It helps local governments, emergency agencies, and residents to plan well in advance to minimize the losses and dangers caused by impending floods.
 Acknowledgments
 This work is part of the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) project, which is funded by the National Science Foundation under award number 2118329.
+Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views of funding sources.
 Appendix A Data and Processing
@@
 If all values in the input tensor at that time steps are equal to $mask\_value$, then the time steps will be masked (skipped) in all downstream layers.
-(a) Flood Evaluator
-(b) Flood Manager
-Figure 9: Flood Evaluator and Flood Manager in FIDLAR.
 Appendix B Model Architecture and Training Hyper-parameter
@@
   • GNN: We consider each time series as one node and the connection as the edge.
-For the variables in Fig.
-6, we assign “1” to the edge between two connected nodes, otherwise, “0”.
+For the variables in Figure 5, we assign “1” to the edge between two connected nodes, otherwise, “0”.
 For the architecture, we stacked two graph convolutional networks (GCNs) with $32$ and $16$ neurons.
@@
 MaxPooling1D layer with pool size $=2$ is added after CNN1D layer.
-  • Transformer: The standard Transformer encoder Vaswani et al.
-[2017] is used.
+  • Transformer: The standard Transformer encoder (Vaswani et al.
+2017) is used.
 In our case, we only used 1 Encoder with 1 head.
@@
 Dropout is used after each layer above.
-  • GTN: We combine the components of GNNs, LSTMs, Transformer, CNNs, and attention, as described in Fig.
-5.
+  • GTN: We combine the components of GNNs, LSTMs, Transformer, CNNs, and attention, as described in Figure 6.
 We have $32$ and $16$ for GNN’s channels, $16$ for LSTM units, $96$ for CNN filters, and one Transformer encoder with $3$ heads (head size=192, embedding dimension=96).
@@
 (8).
-Fig.
-10 provides the corresponding visualization.
+Figure 9 provides the corresponding visualization.
 $x=min\{max\{x,MIN\},MAX\},$ (8)
@@
 In the following table, we show some hyper-parameters used in our experiments: learning rate, batch size, epoch, decay rate, decay step, patience for early stopping, dropout rate, and regularization factors $L_{1}$ and $L_{2}$.
+Table 6: Best hyperparameters of each DL model for Flood Evaluator.
 Methods LR Batch Size Epoch Decay Rate Decay Step Patience Dropout Rate $L_{1}$ $L_{2}$
@@
 RNN 1e-4 512 3000 0.95 10,000 500 0.2 1e-5 1e-5
-CNN 1e-4 512 3000 0.95 10,000 500 0.1 - -
-GCN 5e-4 512 3000 0.95 10,000 500 0.0 - -
-TCN 5e-4 512 3000 0.95 10,000 500 0.1 - -
-RCNN 1e-4 512 3000 0.95 10,000 500 0.2 0 1e-5
-Transformer 5e-4 512 3000 0.95 10,000 500 0.3 - -
+CNN 1e-4 512 3000 0.95 10,000 500 0.1 0.0 0.0
+GCN 5e-4 512 3000 0.95 10,000 500 0.0 0.0 0.0
+TCN 5e-4 512 3000 0.95 10,000 500 0.1 0.0 0.0
+RCNN 1e-4 512 3000 0.95 10,000 500 0.2 0.0 1e-5
+Transformer 5e-4 512 3000 0.95 10,000 500 0.3 0.0 0.0
 GTN (ours) 5e-4 512 3000 0.95 10,000 500 0.5 1e-5 1e-5
-Table 6: Best hyperparameters of each DL model for Flood Evaluator.
+Table 7: Best hyperparameters of each DL model for Flood Manager.
 Methods LR Batch Size Epoch Decay Rate Decay Step Dropout Rate Patience $L_{1}$ $L_{2}$
@@
 GTN (ours) 3e-3 512 700 0.95 10,000 0.5 100 1e-5 1e-5
-Table 7: Best hyperparameters of each DL model for Flood Manager.
 Appendix C More experimental results
 C.1 Flood prediction at all locations
-In Table 2, we provided the comparison of the performance of different models for the Flood Evaluator at time point $t+1$, but at one location (S1).
+In Table 2, we provided the comparison of the performance of models for the Flood Evaluator at time point $t+1$, but at one location (S1).
 In Table 8, we provide the comparison of the performance of different models for the Flood Evaluator on the test set at time $t+1$, but for all locations.
+Our GTN model still provides the best performance in flood prediction.
+Table 8: Comparison of different models for the Flood Evaluator on the test set (at time t+1 for all stations).
+“Over Timesteps” (and “Under Timesteps”) represent the number of time steps during which the water levels exceed the upper threshold (subceed the lower threshold, resp.).
+Similarly, “Over Area” (and “Under Area”) refer to the area between the water level curve and the upper threshold (lower threshold, resp.), as shown in Figure 4 in the manuscript.
 Methods MAE (ft) RMSE (ft) Over Timesteps Over Area Under Timesteps Under Area
@@
 GTN (ours) 0.047 0.064 439 76.01 5,859 1689.45
-Table 8: Comparison of different models for the Flood Evaluator on the test set (at time t+1 for all stations).
-“Over Timesteps” (and “Under Timesteps”) represent the number of time steps during which the water levels exceed the upper threshold (subceed the lower threshold, resp.).
-Similarly, “Over Area” (and “Under Area”) refer to the area between the water level curve and the upper threshold (lower threshold, resp.), as shown in Fig.
-4 in the manuscript.
 C.2 Flood mitigation at all locations
@@
 We provide the comparison results of different models for the Flood Manager on the test set (at time t+1 for all locations) in Table 9.
+All DL-based methods show better results than “rule-based” and GA-based approaches for flood mitigation.
+Our GTN model provides the best performance in flood mitigation and comparable performance in avoiding water wastage.
+Table 9: Comparison of different methods for the Flood Manager on the test set (at time t+1 for all stations).
+The denotes that the GA method was used for Flood Manager with a physics-based (HEC-RAS) as an Evaluator.
+-- denotes too long to get results.
+The denotes that the GA method was used for Flood Manager with the DL-based GTN as the Evaluator.
+All other rows are DL-based flood managers with a DL-based GTN as the Evaluator.
+We mark the best results in bold.
 Method Manager Over Timesteps Over Area Under Timesteps Under Area
 Rule-based - 427 68.61 5,116 1465.10
-GA-based Genetic $*{}^{*}$ (GA) - - - -
-Genetic ${}^{{\dagger}}$ (GA) 343 67.50 1,817 416
+GA-based Genetic Algorithm - - - -
+Genetic Algorithm 343 67.50 1,817 416
 DL-based MLP 435 67.28 3,511 818.06
@@
 Transformer 374 54.31 4,378 1,142.98
-${}^{{\ddagger}}$ (Ours) 114 12.09 617 98.42
-Table 9: Comparison of different methods for the Flood Manager on the test set (at time t+1 for all stations).
-The $*{}^{*}$ denotes that the GA method was used for Flood Manager with a physics-based (HEC-RAS) as an Evaluator.
--- denotes too long to get results.
-The ${}^{{\dagger}}$ denotes that the GA method was used for Flood Manager with the DL-based GTN as the Evaluator.
-All other rows are DL-based flood managers with a DL-based GTN as the Evaluator.
-${}^{{\ddagger}}$ denotes the best.
+GTN (Ours) 114 12.09 617 98.42
 C.3 Flood mitigation with one event
-In Fig.
-7, we visualize the results of different models for the Flood Manager on a short event spanning 18 hours from September 3rd (09:00) to September 4th (03:00) in 2019 for one location of interest.
+In Figure 7, we visualize the results of different models for the Flood Manager on a short event spanning 18 hours from September 3rd (09:00) to September 4th (03:00) in 2019 for one location of interest.
 Below are the corresponding experiment results.
+Our GTN shows the best performance in flood mitigation.
+Table 10: Comparison of different Managers for flood Mitigation (at time t+1 for one event S1). The best results are in bold.
 Method Manager Over Timesteps Over Area Under Timesteps Under Area
 Rule-based - 6 0.866 0 0
-GA-Based Genetic $*{}^{*}$ (GA) 4 0.351 0 0
-Genetic ${}^{{\dagger}}$ (GA) 6 0.764 0 0
+GA-Based Genetic Algorithm (GA) 4 0.351 0 0
+Genetic Algorithm (GA) 6 0.764 0 0
 DL-Based MLP 6 0.614 0 0
@@
 Transformer 6 0.614 0 0
-${}^{{\ddagger}}$ (Ours) 1 0.022 0 0
-Table 10: Comparison of different methods of Flood Manager for flood Mitigation (at time t+1 for one event S1).
+GTN (Ours) 1 0.022 0 0
 Appendix D Visualization
-Fig. 11 presents the results for all locations compared to Fig. 7 for one location of interest. Table 11 shows the corresponding performance measures.
+Figure 10 presents the results for all locations compared to Figure 7 for one location of interest.
+Table 11 shows the corresponding performance measures.
+Our GTN shows the best performance in flood mitigation.
 Figure 10: Visualization for flood mitigation at all locations. Two gray dashed lines represent the upper (3.5 ft) and lower threshold (0.0 ft).
+Table 11: Comparison of different methods of Flood Manager for flood Mitigation (at time t+1 for one event at all locations). We mark the best results in bold.
 Method Manager Over Timesteps Over Area Under Timesteps Under Area
 Rule-based - 23 3.62 0 0
-GA-Based Genetic $*{}^{*}$ (GA) 16 1.39 0 0
-Genetic ${}^{{\dagger}}$ (GA) 23 3.75 0 0
+GA-Based Genetic Algorithm (GA) 16 1.39 0 0
+Genetic Algorithm (GA) 23 3.75 0 0
 DL-Based MLP 23 4.18 0 0
@@
 Transformer 22 3.00 0 0
-${}^{{\ddagger}}$ (Ours) 5 0.22 0 0
-Table 11: Comparison of different methods of Flood Manager for flood Mitigation (at time t+1 for one event at all locations).
+GTN (Ours) 5 0.22 0 0